### v1

# An Efficient Finite Difference Approximation via a Double Sample-Recycling Approach

Abstract
Estimating stochastic gradients is pivotal in fields like service systems within operations research. The classical method for this estimation is the finite difference approximation, which entails generating samples at perturbed inputs. Nonetheless, practical challenges persist in determining the perturbation and obtaining an optimal finite difference estimator in the sense of possessing the smallest mean squared error (MSE). To tackle this problem, we propose a double sample-recycling approach in this paper. Firstly, pilot samples are recycled to estimate the optimal perturbation. Secondly, recycling these pilot samples again and generating new samples at the estimated perturbation, lead to an efficient finite difference estimator. We analyze its bias, variance and MSE. Our analyses demonstrate a reduction in asymptotic variance, and in some cases, a decrease in asymptotic bias, compared to the optimal finite difference estimator. Therefore, our proposed estimator consistently coincides with, or even outperforms the optimal finite difference estimator. In numerical experiments, we apply the estimator in several examples, and numerical results demonstrate its robustness, as well as coincidence with the theory presented, especially in the case of small sample sizes.
Key words: Finite difference; Sample-recycling; Mean square error; Perturbation; Bootstrap.
1 Introduction
Gradient estimation involves estimating a function’s derivative. In realistic scenarios, the exact function expression is often unknown, and one can only approximate it through stochastic simulation. Stochastic gradient estimation finds significant applications in stochastic optimization, machine learning, finance, and operations research.
Stochastic gradient estimation has substantial applications in queuing theory, one of the most important fields of study in operations research (see Fu, 2006; Lam et al., 2022; Li and Lam, 2020). In this field, the primary objective function pertains to the performance of a queuing system, such as average waiting time or queue length. A stochastic service system can be designed and controlled more effectively by analyzing the probabilities involved in queueing and waiting phenomena. Queuing theory finds extensive applications in resource allocation management for manufacturing, product production, transportation, and inventory. Stochastic gradient estimation methods can be employed to estimate the gradient of the queuing model performance function to the model parameters. In this manner, efficient optimization of queueing systems can be achieved, thereby improving overall performance and enhancing the efficiency of operations.
The finite difference technique (see Asmussen and Glynn, 2007; L’Ecuyer, ; Zazanis and Suri, 1993) is a classic approach for estimating the gradient of a function. Forward, backward, and central finite difference (CFD) methods are examples of the finite difference methods. Finite difference estimation necessitates the creation of samples with varying input parameters. The perturbation size is determined by reducing the estimator’s mean squared error (MSE). In contrast, increasing the perturbation leads to both an increase in the bias and a decrease in the variance. To reduce the MSE, the bias and variance must be balanced. Without using common random numbers (CRNs), the optimal perturbation size of the CFD technique is $cn^{-1/6}$, corresponding to the optimal MSE of order $n^{-2/3}$, where $n$ is the number of different pairs in the simulation (Fox and Glynn, 1989). The ideal perturbation size in the forward or backward finite difference method is $c_{1}n^{-1/4}$, and the corresponding optimal MSE order is $n^{-1/2}$. Despite the fact that the choice of perturbation size is widely regarded as the heart of the finite difference technique in relevant studies, the perturbation selection approach has not been fully established. Although the order of the optimal perturbation corresponding to the various finite difference methods is known, the frontal constants, i.e., $c$ and $c_{1}$, remain unknown. Experiments demonstrate that changing the constant value has a significant influence on the MSE, especially for finite samples. However, estimating unknown constants is sometimes seen as more difficult since it requires more model information, such as higher-order derivatives and noise variance. These details are unknown, and even approximating them is a more challenging task than the gradient estimation itself. This work explores the CFD technique, with the goal of extracting this information and providing a more accurate CFD estimator.
The closest work to ours is Li and Lam (2020). They suggest a two-step technique that uses a few samples in the first stage to roughly estimate the third-order derivative of the function using regression and estimate the noise using the sample variance. After this, they obtained an estimator of the optimal perturbation. Then, they construct a finite difference estimator according to the estimated perturbation in the second stage. Li and Lam (2020) referred to their method as the estimation-minimization CFD (EM-CFD) algorithm. In order to more accurately estimate the unknown constants and effectively utilize the samples, we propose a double sample-recycling (DSR) approach to the finite difference estimator for stochastic gradient estimation in this paper. Our work also employs the two-stage strategy but from a sample-recycling angle. That is, we apply two sample-recycling techniques in the two stages to obtain accurate estimators of perturbation and finite difference estimator. Specifically, we make the following contributions.
  1. 1.
    We propose a sample-recycling method based on the bootstrap technique to estimate the constants. The first stage devotes to estimating the constants, primarily relying on the functional relationships between the bias and variance of the finite difference estimator and the perturbation. We apply the sample-recycling technique, specifically the bootstrap method, on a set of pilot samples to estimate the bias and variance. Then the constants can be estimated through regression, utilizing the established functional relationships. The sample-recycling technique not only saves samples but also effectively estimates the bias and variance, so leading to relatively accurate estimators of the constants and then the optimal perturbation.
  2. 2.
    We provide theoretical analyses of the asymptotic properties in the first stage. We show that the bias and variance estimated using the bootstrap method have the same asymptotic properties as the theoretical ones, thereby preserving the aforementioned functional relationships. This ensures the validity of our proposed estimators for the constants and the optimal perturbation. Furthermore, we provide the convergence rate of these estimators, serving as a guide for the implementation of the first stage.
  3. 3.
    We construct a DSR finite difference estimator and prove its theoretical properties In the second stage, we generate remaining samples at the estimated optimal perturbation. Then we combine these samples with the pilot ones to construct the finite difference estimator. Specifically, we recycle the pilot samples by transforming them to ones with the same bias and variance as the remaining samples. Our theoretical analysis reveals a reduction in the variance of the DSR estimator compared to the optimal finite difference estimator and, in some cases, a decrease in bias.
  4. 4.
    We propose an efficient algorithm for practical implementation. Theoretical properties provide a guiding framework for practical implementation. Based on this, we propose an efficient algorithm to derive a robust finite difference estimator in practice. In particular, we apply the weighted least squares to mitigate the heteroscedasticity in the regression.
  5. 5.
    We demonstrate the practical benefits of DSR method via numerical experiments. We apply the DSR method in several examples. The results obtained from these numerical examples demonstrate the good performance of the proposed DSR method in terms of estimators of the constants, perturbation, and the finite difference estimator. The numerical results indicate the advantage of the DSR method over other approaches, especially in the case of small sample sizes.
The rest of this paper is organized as follows. In Section 2, we give some background on how to use the CFD method for gradient estimation. In Section 3, we introduce the bootstrap method, apply it to gradient estimation and provide theoretical guarantees for it. Section 4 guides the process of reusing samples and provides theoretical evidence for the effectiveness of this operation. For implementation, an efficient algorithm is proposed in Section 5 and in Section 6, we give some experiments to verify the theoretical results, followed by conclusions in Section 7. Proofs are provided in the appendix.
2 Background
In this section, we introduce the background of the finite difference method for stochastic gradient estimation. More details can be found in Asmussen and Glynn (2007), Fu (2006) and Glasserman (2013). Consider the models that depend on a single parameter $\theta$, where $\theta$ varies within some range $\Theta\subset\mathbb{R}$. Denote that $\alpha(\cdot)$ is a performance measure of interest and assume that we can only estimate it by simulation. Within the simulation for any chosen $\theta\in\Theta$, each trail gives an unbiased but noisy estimate of $\alpha(\theta)$, denoted by $Y(\theta)$, i.e., $\alpha(\theta)=\mathbb{E}[Y(\theta)]$. Suppose that we do not use the CRNs, i.e., for $\theta_{1}\neq\theta_{2}\in\Theta$, $Y(\theta_{1})$ and $Y(\theta_{2})$ are independent. We would like to estimate the first-order derivative $\alpha^{\prime}(\theta_{0})$, where $\theta_{0}\in\Theta$ is the point of interest. To estimate $\alpha^{\prime}(\theta_{0})$, we prefer the CFD method over the forward (or backward) finite difference method because compared to the forward (and backward) finite difference method, the CFD method yields an improvement in the convergence rate of the bias. It is worth mentioning that the methods in Sections 3 and 4 are seamlessly adapted to the forward (and backward) finite difference scheme.
The CFD scheme is similar to the definition of the derivative and utilizes the information at the neighboring points on both sides of $\theta_{0}$. That is, $\alpha^{\prime}(\theta)$ is approximated by
$\widetilde{\alpha}_{h}^{\prime}(\theta_{0})=\frac{\alpha(\theta_{0}+h)-\alpha(\theta_{0}-h)}{2h},$ (1)
where $h$ is a perturbation parameter. As $h$ tends to 0, $\widetilde{\alpha}_{h}^{\prime}(\theta_{0})$ tends to $\alpha^{\prime}(\theta_{0})$. Denote
$\Delta(h)=\frac{Y(\theta_{0}+h)-Y(\theta_{0}-h)}{2h}.$ (2)
Evidently, $\Delta(h)$ is a noisy estimator of $\alpha^{\prime}(\theta_{0})$ and this estimator is the output of the CFD method.
Specifically, the CFD scheme specifies a perturbation parameter $h>0$. At the point $\theta_{0}+h$, the simulation is repeated independently, and $n$ independently and identically distributed (i.i.d.) observations $\{Y_{1}(\theta_{0}+h),...,Y_{n}(\theta_{0}+h)\}$ are obtained. Likewise, at $\theta_{0}-h$, $n$ i.i.d. observations $\{Y_{1}(\theta_{0}-h),...,Y_{n}(\theta_{0}-h)\}$ are generated. Using the sample average, we construct a CFD estimator of $\alpha^{\prime}(\theta_{0})$, denoted by
$\widehat{\Delta}_{n,h}\triangleq\frac{1}{n}\sum_{i=1}^{n}\Delta_{i}(h)=\frac{1}{n}\sum_{i=1}^{n}\frac{Y_{i}(\theta_{0}+h)-Y_{i}(\theta_{0}-h)}{2h}.$ (3)
For any $i=1,...,n$, $Y_{i}(\theta_{0}+h)$ and $Y_{i}(\theta_{0}-h)$ are the unbiased estimators of $\alpha(\theta_{0}+h)$ and $\alpha(\theta_{0}-h)$, respectively, so $\Delta_{i}(h)$ and $\widehat{\Delta}_{n,h}$ are both the unbiased estimators of $\widetilde{\alpha}_{h}^{\prime}(\theta_{0})$. Let $Z_{i}(h)$ denote the zero-mean error associated with $\Delta_{i}(h)$, i.e., $\Delta_{i}(h)=\widetilde{\alpha}_{h}^{\prime}(\theta_{0})+Z_{i}(h)$. Then, denote
$\bar{Z}_{n,h}=\frac{1}{n}\sum_{i=1}^{n}Z_{i}(h)$
be the zero-mean error associated with $\widehat{\Delta}_{n,h}$, i.e.,
$\widehat{\Delta}_{n,h}=\widetilde{\alpha}_{h}^{\prime}(\theta_{0})+\bar{Z}_{n,h}.$ (4)
The selection of $h$ holds significant importance as it directly influences the quality of $\widehat{\Delta}_{n,h}$. If $h$ is too large, the expected value of $\widehat{\Delta}_{n,h}$, i.e., $\widetilde{\alpha}_{h}^{\prime}(\theta_{0})$ could deviate significantly from $\alpha^{\prime}(\theta_{0})$, resulting in an unacceptable bias. Conversely, if $h$ is too small, the variance of $\widehat{\Delta}_{n,h}$ will explode due to its position in the denominator. To choose a proper $h$, the challenge lies in striking a balance between mitigating the bias and restraining the variance of $\widehat{\Delta}_{n,h}$. To calculate the bias and variance elaborately, we make the following assumptions.
Assumption 1.
$\alpha(\theta)$ is thrice continuously differentiable in a neighborhood of $\theta_{0}$ and $\alpha^{(3)}(\theta_{0})\neq 0$.
Assumption 2.
The standard deviation of $Y(\theta)$, denoted by $\sigma(\theta)$, is continuous at $\theta_{0}$ and $\sigma(\theta)$ is bounded below away from 0.
Assumptions 1 and 2 can be found in many other works, such as Glasserman (2013). Under Assumption 1, according to the Taylor expansion, we have
$\alpha(\theta_{0}+h)=\alpha(\theta_{0})+\alpha^{\prime}(\theta_{0})h+\frac{\alpha^{(2)}(\theta_{0})}{2}{h}^{2}+\frac{\alpha^{(3)}(\theta_{0})}{6}{h}^{3}+o(h^{3}),$
$\alpha(\theta_{0}-h)=\alpha(\theta_{0})-\alpha^{\prime}(\theta_{0})h+\frac{\alpha^{(2)}(\theta_{0})}{2}h^{2}-\frac{\alpha^{(3)}(\theta_{0})}{6}h^{3}+o(h^{3}).$
Subtraction eliminates $\alpha^{(2)}(\theta_{0})$, leaving
$\widetilde{\alpha}_{h}^{\prime}(\theta_{0})=\alpha^{\prime}(\theta_{0})+Bh^{2}+o(h^{2}),$
where $B=\alpha^{(3)}(\theta_{0})/6$. Therefore, the bias of $\widehat{\Delta}_{n,h}$ is
$\mathrm{Bias}\left[\widehat{\Delta}_{n,h}\right]=\mathbb{E}\widehat{\Delta}_{n,h}-\alpha^{\prime}(\theta_{0})=\widetilde{\alpha}_{h}^{\prime}(\theta_{0})-\alpha^{\prime}(\theta_{0})=Bh^{2}+o(h^{2}).$ (5)
From (3), the variance of $\widehat{\Delta}_{n,h}$ is
$\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]=\frac{1}{n}\mathrm{Var}\left[\frac{Y(\theta_{0}+h)+Y(\theta_{0}-h)}{2h}\right]=\frac{\sigma^{2}(\theta_{0}+h)+\sigma^{2}(\theta_{0}-h)}{4nh^{2}}.$
Under Assumption 2, we have $\sigma^{2}(\theta_{0}+h)=\sigma^{2}(\theta_{0})+o(1)$ and $\sigma^{2}(\theta_{0}-h)=\sigma^{2}(\theta_{0})+o(1)$. Therefore,
$\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]=\frac{\sigma^{2}(\theta_{0})+o(1)}{2nh^{2}}.$ (6)
Then the MSE of $\widehat{\Delta}_{n,h}$ is
$\mbox{MSE}\left[\widehat{\Delta}_{n,h}\right]=\mathrm{Bias}^{2}\left[\widehat{\Delta}_{n,h}\right]+\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]=\left(B+o(1)\right)^{2}h^{4}+\frac{\sigma^{2}(\theta_{0})+o(1)}{2nh^{2}}.$ (7)
It follows from (7) that $h$ controls a trade-off between the bias and variance, and thus we should choose an appropriate $h$ to minimize the MSE. According to the relationship between the arithmetic average and the geometric average, we get
$B^{2}h^{4}+\frac{\sigma^{2}(\theta_{0})}{2nh^{2}}\geq 3\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/3},$
where the equality holds if and only if
$h^{*}=\left(\frac{\sigma^{2}(\theta_{0})}{4nB^{2}}\right)^{1/6}.$ (8)
The optimal bias, variance and MSE are thus
$\begin{cases}\mbox{Bias}^{*}=\left(\dfrac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+o\left(n^{-1/3}\right),\\ \mathrm{Var}^{*}=\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{2n^{2}}\right)^{1/3}+o\left(n^{-2/3}\right),\\ \mbox{MSE}^{*}=3\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/3}+o\left(n^{-2/3}\right).\end{cases}$ (9)
In order to obtain an estimator $\widehat{\Delta}_{n,h}$ with optimal MSE, it is important to determine the value of $h^{*}$ according to (8). Although it is established that $h^{*}=Cn^{-1/6}$ from (8), where $C=\left[\sigma^{2}(\theta_{0})/(4B^{2})\right]^{1/6}$, the constant $C$ is typically unknown, and involves two constants $\alpha^{(3)}(\theta_{0})$ and $\sigma^{2}(\theta_{0})$ that depend on model information. However, estimating these two constants is more challenging than estimating the gradient itself because they involve the higher-order derivatives of $\alpha(\theta)$. Furthermore, experimental findings indicate that assigning a perturbation with a significant deviation from the true one leads to substantial discrepancies in the MSE of the CFD estimator. These challenges motivate us to explore an efficient approach to estimate the constant $C$ so that the CFD estimator achieves its optimal MSE.
3 Sample Recycling for Perturbation Selection
In this section, we propose a sample recycling procedure aimed at estimating $C$, specifically $B$ and $\sigma^{2}(\theta_{0})$. Our inspiration stems from equations (5) and (6), which reveal that $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ are linear w.r.t. $h^{2}$ and $1/h^{2}$, respectively. If we can accurately calculate $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ for any given $h$, then choosing $h_{1}\neq h_{2}$ and solving the system of linear equations
$\begin{cases}\mathbb{E}\widehat{\Delta}_{n,h_{1}}-\alpha^{\prime}(\theta_{0})=Bh_{1}^{2}+o(h_{1}^{2}),\\ \mathbb{E}\widehat{\Delta}_{n,h_{2}}-\alpha^{\prime}(\theta_{0})=Bh_{2}^{2}+o(h_{2}^{2}),\end{cases}$ (10)
leads to an estimator of $B$. In this case, the estimator of $B$ has zero variance, and its bias tends to 0 as $h_{1},h_{2}\to 0$ because the vanishing rate of $o(h_{1}^{2})$ and $o(h_{2}^{2})$ surpasses that of $h_{1}^{2}$ and $h_{2}^{2}$, respectively. Likewise, $\sigma^{2}(\theta_{0})$ can be estimated by $2nh_{1}^{2}\mathrm{Var}\left[\widehat{\Delta}_{n,h_{1}}\right]$.
However, $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ are typically unknown, and estimating them may be as challenging as estimating $\alpha^{\prime}(\theta_{0})$. Therefore, we turn to seek their surrogates. A straightforward approach is to use sample mean and variance, respectively. That is, we generate several sample versions of $\widehat{\Delta}_{n,h}$, denoted by $\left\{\widehat{\Delta}_{n,h}(l),l=1,2,...,p\right\}$, and then use their sample mean and variance as the surrogates for $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$, respectively. Nevertheless, for any perturbation $h$, we need $np$ sample pairs [^4] Note that sometimes we also use “samples”. Throughout, the meanings of “samples” and “sample pairs” are same and we do not make distinctions. to obtain these surrogates, which is unavailable due to the limitation of sample size, especially when $n$ or $p$ is large. To address this limitation, we introduce a sample recycling technique, the bootstrap, to construct the sample mean and variance.
On the other hand, our primary objective is to estimate $\alpha^{\prime}(\theta_{0})$ rather than determining the optimal perturbation $h^{*}$. Consequently, the number of samples required for estimating $h^{*}$ should be relatively small compared to the total number of samples available. As a result, using bootstrap sample means as surrogates of the expectations in (10) may not provide sufficient accuracy to yield a relatively precise estimator of $B$ by solving the two equations in (10). To address this issue, we select $K$ different values of $h$, expanding (10) to encompass $K$ linear equations correspondingly. Then least-squares regression is applied to derive the slope of linear equations, i.e., the constant $B$.Similar procedure is also used to obtain the estimator of $\sigma^{2}(\theta_{0})$.
In the following, we will introduce the bootstrap technique for estimating the sample mean and variance and the least-squares regression procedure for estimating the constants, along with their theoretical guarantees.
3.1 Bootstrap sample mean and variance
In this subsection, we apply the bootstrap technique to establish surrogates of $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ based on some pilot samples. We also provide their corresponding theoretical analyses.
The bootstrap technique is described as follows. For a fixed $h$, assume that $n_{b}$ sample pairs are generated, denoted by $\{(Y_{i}(\theta_{0}+h),Y_{i}(\theta_{0}-h)),i=1,2,...,n_{b}\}$. Correspondingly, as defined in (2), we obtain $n_{b}$ outputs $\bm{\Delta}(h)\triangleq\{\Delta_{i}(h),i=1,2,...,n_{b}\}$. We then select, with replacement, $n_{b}$ times from $\bm{\Delta}(h)$ independently and randomly to obtain a group of bootstrap samples $\bm{\Delta}^{*}(h)=\{\Delta_{i}^{*}(h),i=1,2,...,n_{b}\}$. Using these bootstrap samples, the estimator of $\alpha^{\prime}(\theta_{0})$ is expressed as
$\widehat{\Delta}^{b}_{n_{b},h}=\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}\Delta_{i}^{*}(h).$
It is worth mentioning that in this procedure, the bootstrap samples come from the same pilot samples (the $n_{b}$ sample pairs) and do not require the generation of extra samples. In essence, the bootstrap technique is based on the reuse of the pilot samples.
It is natural to expect that the bootstrap mean and variance of $\widehat{\Delta}^{b}_{n_{b},h}$ will exhibit similar asymptotic properties as described in (5) and (6), respectively, when $n_{b}$ is sufficiently large. However, establishing these properties theoretically is not a straightforward task, because $\widehat{\Delta}^{b}_{n_{b},h}$ is derived from the resampling technique given the pilot samples. Therefore, these properties should be established based on the given pilot samples.
For the purposes described above, let $\mathbb{E}_{*}$ and $\mathrm{Var}_{*}$ denote the expectation and variance under the bootstrap probability measure $\mathbb{P}_{*}(\cdot)\triangleq\mathbb{P}(\cdot|\bm{\Delta}(h))$. Note that $\mathbb{P}_{*}$ is a probability conditional on pilot samples. Also, $\mathbb{P}_{*}$ is a discrete probability distribution, and its probability mass function is shown in Table 1.
Table 1: The value and probability of $\mathbb{P}_{*}$.
Value $\Delta_{1}(h)$ $\Delta_{2}(h)$cdots $\Delta_{n_{b}}(h)$
Probability $1/n_{b}$ $1/n_{b}$cdots $1/n_{b}$
The asymptotic properties of the bootstrap mean and variance of $\widehat{\Delta}^{b}_{n_{b},h}$ are summarized in Theorem 1. Before proceeding, we make the following Assumption 3, which facilitates the investigation of the term $o(h^{2})$ in (5). Furthermore, this assumption aids in the analysis of the estimation error of $B$, which in turn assists in setting the initial perturbations.
Assumption 3.
$\alpha(\theta)$ is five times continuously differentiable in a neighborhood of $\theta_{0}$ and $\alpha^{(5)}(\theta_{0})\neq 0$.
In many practical problems, the function $\alpha(\theta)$ is five times continuously differentiable. For example, in the $M/M/1$ queue, the average waiting time is a smooth function of the arrival and service rate (see Kleinrock and Gail, 1996). Now we provide the theorem, and the proof is provided in Appendix A.1.
Theorem 1.
Denote $\nu_{4}(h)=\mathbb{E}\left[Z_{1}(h)\right]^{4}$ and $\nu_{4}=\underset{h\to 0}{\lim}\mathbb{E}\left[hZ_{1}(h)\right]^{4}$. If $\nu_{4}<\infty$, then under Assumptions 2 and 3, as $n_{b}\to\infty$,
$\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h}=\alpha^{\prime}(\theta_{0})+Bh^{2}+Dh^{4}+o(h^{4})+\bar{Z}_{n_{b},h},$ (11)
$\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]=\frac{(n_{b}-1)(\sigma^{2}(\theta_{0})+o(1))}{2n_{b}^{2}h^{2}}+\phi(h),$ (12)
where $D=\alpha^{(5)}(\theta_{0})/120$ and $\phi(h)$ is a zero-mean error with
$\mathrm{Var}[\phi(h)]=\frac{(n_{b}-1)^{2}}{n_{b}^{4}h^{4}}\left(\frac{\nu_{4}+o(1)}{n_{b}}-\frac{n_{b}-3}{n_{b}(n_{b}-1)}\frac{\sigma^{4}(\theta_{0})+o(1)}{4}\right).$ (13)
Theorem 1 indicates that by the resampling technique, we obtain the bootstrap samples whose mean and variance under $\mathbb{P}_{*}$ exhibit similar properties to those described in (5) and (6), respectively. However, it is worth mentioning the difference between Theorem 1 and (5) and (6). In particular, (11) eliminates the randomness from the bootstrap sampling in $\{\Delta_{1}(h),\Delta_{2}(h),...,\Delta_{n_{b}}(h)\}$, and (12) measures the degree of dispersion under $\mathbb{P}_{*}$. In contrast, (5) eliminates the randomness of simulating $\widehat{\Delta}_{n,h}$, and (6) measures the degree of dispersion of $\widehat{\Delta}_{n,h}$.
Although $n_{b}$ is the number of pilot samples and so it could be small compared with the number of total samples, the asymptotic results in Theorem 1 serve as a guide for illustrating the properties of $\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h}$ and $\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]$. Specifically, Theorem 1 demonstrates that $\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h}$ and $\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]$ inherit the asymptotic properties of $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$, respectively. That is, they are linear w.r.t. $h^{2}$ and $1/h^{2}$, respectively. As a result, we can estimate $B$ and $\sigma^{2}(\theta_{0})$ through these linear relationships, which we will present in the next subsection.
3.2 Regression for perturbation selection
As mentioned at the beginning of Section 3, to obtain relatively precise estimators of $B$ and $\sigma^{2}(\theta_{0})$, multiple equations like (10) have to be constructed, and least-squares regression is used to derive the slopes of these linear equations, i.e., the constants $B$ and $\sigma^{2}(\theta_{0})$.
In our setting, we consider $K$ perturbation parameters $h_{1},h_{2},...,h_{K}$, where $K$ is finite. Using $Kn_{b}$ pilot samples, we construct a system of $K$ linear equations like (10) by (11) in Theorem 1, and we consider the vector version as follows.
$\bm{Y}_{e}$ $=\bm{X}_{e}\bm{\beta}_{e}+\mathcal{E}_{e},$ (14)
where
$\quad\quad\quad\bm{Y}_{e}=\left[\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h_{1}},...,\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h_{K}}\right]^{\top},\bm{X}_{e}=\left[\begin{array}[]{lll}1&...&1\\ h_{1}^{2}&...&h_{K}^{2}\end{array}\right]^{\top},$ (15)
$\bm{\beta}_{e}=\left[\alpha^{\prime}(\theta_{0}),B\right]^{\top},\mathcal{E}_{e}=\left[Dh_{1}^{4}+o(h_{1}^{4})+\bar{Z}_{n_{b},h_{1}},...,Dh_{K}^{4}+o(h_{K}^{4})+\bar{Z}_{n_{b},h_{K}}\right]^{\top}.$
Similarly, by (12) in Theorem 1, the system of $K$ linear equations for estimating $\sigma^{2}(\theta_{0})$ is
$\bm{Y}_{v}$ $=\bm{X}_{v}\bm{\beta}_{v}+\mathcal{E}_{v},$ (16)
where
$\bm{Y}_{v}=\left[\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h_{1}}\right],...,\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h_{K}}\right]\right]^{\top},\ \bm{X}_{v}=\left[\frac{n_{b}-1}{2n_{b}^{2}h_{1}^{2}},...,\frac{n_{b}-1}{2n_{b}^{2}h_{K}^{2}}\right]^{\top},$ (17)
$\bm{\beta}_{v}=\sigma^{2}(\theta_{0}),\ \mathcal{E}_{v}=\left[\frac{(n_{b}-1)o(1)}{2n_{b}^{2}h_{1}^{2}}+\phi(h_{1}),...,\frac{(n_{b}-1)o(1)}{2n_{b}^{2}h_{K}^{2}}+\phi(h_{K})\right]^{\top}.$
Using the least-squares method, we derive the estimators of $\bm{\beta}_{e}$ and $\bm{\beta}_{v}$, denoted by $\widehat{\bm{\beta}}_{e}$ and $\widehat{\bm{\beta}}_{v}$, respectively, where
$\widehat{\bm{\beta}}_{e}=\left(\bm{X}_{e}^{\top}\bm{X}_{e}\right)^{-1}\bm{X}_{e}^{\top}\bm{Y}_{e},\quad\widehat{\bm{\beta}}_{v}=\left(\bm{X}_{v}^{\top}\bm{X}_{v}\right)^{-1}\bm{X}_{v}^{\top}\bm{Y}_{v}.$
Let the first term of $\widehat{\bm{\beta}}_{e}$ be $\widehat{\alpha}^{\prime}(\theta_{0})$ and the second term be $\widehat{B}$. In the following, we show that $\widehat{\bm{\beta}}_{e}$ and $\widehat{\bm{\beta}}_{v}$ are consistent estimators of $\bm{\beta}_{e}$ and $\bm{\beta}_{v}$, respectively.
3.2.1 Consistency of $\widehat{B}$
The following theorem presents the consistency of $\widehat{B}$, whose proof is provided in Appendix A.2. The conditions about $h_{k}(k=1,...,K)$ in Theorem 2 ensure that the regression error vanishes fast enough, which is necessary for the consistency of regression estimators. Note that we have assumed $h_{k}=c_{k}n_{b}^{\gamma}$ $(k=1,...,K)$ in the theorem, where $\gamma<0$, and $c_{k}\neq 0$ are constants. This choice of $h_{k}$ aims to minimize the second moment of the regression error $\mathcal{E}_{e}$ in (15).
Theorem 2.
Suppose that Assumptions 1 and 3 hold. For any $k=1,...,K(K\geq 2)$, let $h_{k}=c_{k}n_{b}^{\gamma}$ $(c_{k}\neq 0,\gamma<0)$ and for any $j\neq k$, $c_{j}\neq c_{k}$. Then,
$\mathbb{E}\left[\widehat{B}\right]-B=H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right),\quad\mathrm{Var}\left[\widehat{B}\right]=V_{K}\frac{\sigma^{2}(\theta_{0})+o(1)}{2n_{b}^{1+6\gamma}},$ (18)
where $H_{K}$ and $V_{K}$ are constants depending on $c_{1},...,c_{K}$. Specifically,
$H_{K}=\frac{K\sum_{k=1}^{K}Dc_{k}^{6}-\sum_{k=1}^{K}c_{k}^{2}\sum_{k=1}^{K}Dc_{k}^{4}}{K\sum_{k=1}^{K}c_{k}^{4}-\left(\sum_{k=1}^{K}c_{k}^{2}\right)^{2}},\quad V_{K}=\frac{-K^{2}\sum_{k=1}^{K}c_{k}^{2}+\left(\sum_{k=1}^{K}c_{k}^{2}\right)^{2}\sum_{k=1}^{K}1/c_{k}^{2}}{\left(K\sum_{k=1}^{K}c_{k}^{4}-\left(\sum_{k=1}^{K}c_{k}^{2}\right)^{2}\right)^{2}}.$
In addition, if $-1/6<\gamma<0$, then $\widehat{B}\stackrel{{\scriptstyle p}}{{\longrightarrow}}B$, where $\stackrel{{\scriptstyle p}}{{\longrightarrow}}$ means convergence in probability.
Theorem 2 indicates that for all $h_{k}=c_{k}n_{b}^{\gamma}$, where $k=1,...,K$ and $c_{k}\neq 0$, as long as $-1/6<\gamma<0$, $\widehat{B}$ is a consistent estimator of $B$. Furthermore, from (18), the convergence rate of the bias of $\widehat{B}$ is $n_{b}^{2\gamma}$, and that of the variance of $\widehat{B}$ is $n_{b}^{-1-6\gamma}$. In addition, Theorem 2 provides valuable guidance for the selection of appropriate values of $h_{k}$ ($k=1,...,K$) during the pilot stage.
Except for fixed values of $c_{1},...,c_{K}$, they can be generated from a proper distribution $\mathcal{P}_{0}$. In this case, $H_{K}$ and $V_{K}$ are both random. Since $c_{1},...,c_{K}$ are coefficients in the perturbations, they must not equal 0, and are typically bounded. For these purposes, there are many optional distributions, such as the truncated normal distribution, denoted by $\psi\left(\mu_{0},\sigma_{0}^{2},L,U\right)$. Specifically, for any $k=1,...,K$, let $c_{k}\stackrel{{\scriptstyle d}}{{=}}X\mathbbm{1}_{\{L\leq X\leq U\}}$, where =$\stackrel{{\scriptstyle d}}{{=}}$ represents the same distribution, $X\sim\mathcal{N}\left(\mu_{0},\sigma_{0}^{2}\right)$, and $L>0$ and $U>0$ denote the lower bound and upper bound, respectively. Because $0<L\leq c_{k}\leq U$, $H_{K}$ and $V_{K}$ are bounded.
When $c_{1},...,c_{K}$ are generated from $\mathcal{P}_{0}$, the bias of $\widehat{B}$ is
$\mbox{Bias}\left[\widehat{B}\right]=\mathbb{E}\left[\mathbb{E}\left[\widehat{B}-B\big{|}\mathcal{P}_{0}\right]\right]=\mathbb{E}[H_{K}]n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right),$ (19)
where the last equality is because $c_{1},...,c_{K}$ are independent with the pilot samples and the bias result in Theorem 2. If $c_{k}\sim\psi\left(\mu_{0},\sigma_{0}^{2},L,U\right)$ for $k=1,...,K$, then $c_{k}$’s are bounded. Therefore, in equation (19), $H_{K}$ and the constants in $o\left(n_{b}^{2\gamma}\right)$ are both bounded.
Likewise, the variance of $\widehat{B}$ is
$\mathrm{Var}\left[\widehat{B}\right]$ $=\mathrm{Var}\left[\mathbb{E}\left[\widehat{B}\big{|}\mathcal{P}_{0}\right]\right]+\mathbb{E}\left[\mathrm{Var}\left[\widehat{B}\big{|}\mathcal{P}_{0}\right]\right]$
$=\mathrm{Var}\left[B+H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right)\right]+\mathbb{E}\left[V_{K}\frac{\sigma^{2}(\theta_{0})+o(1)}{2n_{b}^{1+6\gamma}}\right]$
$=\mathrm{Var}[H_{K}]n_{b}^{4\gamma}+o\left(n_{b}^{4\gamma}\right)+\mathbb{E}[V_{K}]\frac{\sigma^{2}(\theta_{0})+o(1)}{2n_{b}^{1+6\gamma}}.$ (20)
Comparing (18) with (19) and (3.2.1), it becomes evident that when $c_{k}$’s are generated from $\mathcal{P}_{0}$, the bias of $\widehat{B}$ remains $O\left(n_{b}^{2\gamma}\right)$, while the variance of $\widehat{B}$ increases by $O\left(n_{b}^{4\gamma}\right)$. However, it is important to note that the convergence rate of the MSE of $\widehat{B}$ remains unchanged. Therefore, during the pilot stage, the $h_{k}$’s selected based on the tradeoff between bias and variance in (18) maintain the same order as those selected using the tradeoff between bias and variance in (19) and (3.2.1).
Note that the first term of $\widehat{\bm{\beta}}_{e}$ is also an estimator of $\alpha^{\prime}(\theta_{0})$, and the second term is $\widehat{B}$. Therefore, similar analysis to the proof of Theorem 2 could yield the asymptotic properties of the first term. If the total simulation budget is used in the pilot stage, then the first term may seem like a reasonable estimator of $\alpha^{\prime}(\theta_{0})$. However, this estimator lacks efficiency, because all perturbations in the pilot stage are not taken at the optimal value.
Similar to Theorem 2, we provide Corollary 6 in Appendix A.3, which concludes the asymptotic property of $\widehat{\alpha}^{\prime}(\theta_{0})$.
3.2.2 Consistency of $\widehat{\beta}$v
To conduct a more precise analysis of the convergence of $\widehat{\bm{\beta}}_{v}$, additional information about $\sigma^{2}(\theta)$ is required. Hence, we provide the following Assumption 4.
Assumption 4.
$\sigma(\theta)$ is continuously differentiable at $\theta_{0}$ and $\sigma^{\prime}(\theta_{0})$ is bounded.
The rationality of Assumption 4 is provided in Remark 1. Under Assumption 4, we are able to refine (12) as follows,
$\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]=\frac{(n_{b}-1)\left[\sigma^{2}(\theta_{0})+\sigma^{\prime}(\theta_{0})^{2}h^{2}+o(h^{2})\right]}{2n_{b}^{2}h^{2}}+\phi(h).$ (21)
The proof of (21) is provided in Appendix A.4. It follows from (21) that $\mathcal{E}_{v}$ in (17) is expressed as
$\mathcal{E}_{v}=\left[\frac{(n_{b}-1)\left[\sigma^{\prime}(\theta_{0})^{2}h_{1}^{2}+o(h_{1}^{2})\right]}{2n_{b}^{2}h_{1}^{2}}+\phi(h_{1}),...,\frac{(n_{b}-1)\left[\sigma^{\prime}(\theta_{0})^{2}h_{K}^{2}+o(h_{K}^{2})\right]}{2n_{b}^{2}h_{K}^{2}}+\phi(h_{K})\right]^{\top}.$ (22)
Evidently, under Assumption 4, the terms $o(1)$’s in (12) and (17) are explicitly expressed, which helps us analyze the bias and variance of $\widehat{\bm{\beta}}_{v}$ more precisely.
Now, we present Theorem 3 which demonstrates the convergence rate of the bias and variance of $\widehat{\bm{\beta}}_{v}$. The proof of Theorem 3 is provided in Appendix A.5.
Theorem 3.
Suppose that Assumptions 2 and 4 hold. For any $k=1,...,K$ $(K\geq 1)$, let $h_{k}=c_{k}n_{b}^{\gamma}$ $(c_{k}\neq 0,\gamma<0)$ and for any $j\neq k$, $c_{j}\neq c_{k}$. Then, as $n_{b}\to\infty$,
$\mathbb{E}\left[\widehat{\bm{\beta}}_{v}\right]-\bm{\beta}_{v}=\widehat{H}_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right),\quad\mathrm{Var}\left[\widehat{\bm{\beta}}_{v}\right]=\widehat{V}_{K}\frac{4\nu_{4}(n_{b}-1)-\sigma^{4}(\theta_{0})(n_{b}-3)}{n_{b}(n_{b}-1)}+o\left(\frac{1}{n_{b}}\right),$
where $\widehat{H}_{K}$ and $\widehat{V}_{K}$ are constants depending on $c_{1},...,c_{K}$. Specifically,
$\widehat{H}_{K}=\sigma^{\prime}(\theta_{0})^{2}\left(\sum_{k=1}^{K}\frac{1}{c_{k}^{4}}\right)^{-1}\sum_{k=1}^{K}\frac{1}{c_{k}^{2}},\quad\widehat{V}_{K}=\left(\sum_{k=1}^{K}\frac{1}{c_{k}^{4}}\right)^{-2}\sum_{k=1}^{K}\frac{1}{c_{k}^{8}}.$
Consequently, we have $\widehat{\bm{\beta}}_{v}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\bm{\beta}_{v}$.
Remark 1.
Assumption 4 serves the purpose of estimating the convergence rate of $\mathrm{Bias}\left[\widehat{\bm{\beta}}_{v}\right]$ and substantiating the appropriateness of selecting $\gamma$. If Assumption 4 is violated, the non-random component of $\mathcal{E}_{v}$ transitions to $o(1)$ (refer to (17)), which could have a slower convergence rate than $O(h^{2})$ but still tends to 0. Meanwhile, the random component remains unaffected. For example, if $\theta_{0}=0$, $\sigma(\theta)=1+\sqrt{\theta}$ when $\theta>0$, $\sigma(\theta)=1-\sqrt{-\theta}$ when $-1/4<\theta\leq 0$ and $\sigma(\theta)=1/2$ when $\theta\leq-1/4$. Then, $\sigma^{\prime}(\theta_{0})$ is not bounded above and for any $40<h\leq 1/$ 4, it holds that $\sigma(\theta_{0}+h)=\sigma(\theta_{0})+\sqrt{h}$, $\sigma(\theta_{0}-h)=\sigma(\theta_{0})-\sqrt{h}$. Additionally, the expression for $\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]$ is given by $(n_{b}-1)(\sigma^{2}(\theta_{0})+h)/(2n_{b}^{2}h^{2})+\phi(h)$. Therefore, for any $h_{1},...,h_{K}\in(0,1/4]$,
$\mathcal{E}_{v}=\left[\frac{(n_{b}-1)h_{1}}{2n_{b}^{2}h_{1}^{2}}+\phi(h_{1}),...,\frac{(n_{b}-1)h_{K}}{2n_{b}^{2}h_{K}^{2}}+\phi(h_{K})\right]^{\top}.$
Consequently, the convergence rate of $\mathrm{Bias}\left[\widehat{\bm{\beta}}_{v}\right]$ is slower than what is illustrated in Theorem 3. At this point, the selection of $\gamma$ may cause confusion. This is because the convergence rate of $\mathrm{Bias}\left[\widehat{\bm{\beta}}_{v}\right]$ is slower than that of $\mathrm{Bias}\left[\widehat{B}\right]$, whereas the convergence rate of $\mathrm{Var}\left[\widehat{\bm{\beta}}_{v}\right]$ is faster than that of $\mathrm{Var}\left[\widehat{B}\right]$. In summary, Assumption 4 provides a greater rationale for the selection of $\gamma$.
Theorem 3 indicates that for all $h_{k}=c_{k}n_{b}^{\gamma}$, where $k=1,...,K$ and $c_{k}\neq 0$, as long as $\gamma<0$, $\widehat{\bm{\beta}}_{v}$ is a consistent estimator of $\bm{\beta}_{v}$. The convergence rate of the variance of $\widehat{\bm{\beta}}_{v}$ is $n_{b}^{-1}$, which is independent of $\gamma$ and faster than that of $\widehat{B}$. Additionally, the convergence rate of the bias of $\widehat{\bm{\beta}}_{v}$ is $n_{b}^{2\gamma}$, which is equal to that of $\widehat{B}$. As a result, when comparing to $\widehat{B}$, $\widehat{\bm{\beta}}_{v}$ is easier to estimate.
3.3 The optimal perturbation in the pilot stage
From the asymptotic properties of $\widehat{B}$ and $\widehat{\bm{\beta}}_{v}$ as shown in Theorems 2 and 3, respectively, we derive that the optimal perturbation in the pilot stage is not the same as the optimal perturbation $h^{*}$ of CFD problem as shown in (8). In fact, since $\mathrm{Var}\left[\widehat{\bm{\beta}}_{v}\right]$ is unrelated to $\gamma$, we only have to focus on (18) and choose a suitable $\gamma$ to balance the bias and variance of $\widehat{B}$. Specifically, according to Theorem 2, we consider the following optimization problem:
$\underset{\gamma,c_{1},...,c_{K}}{\mbox{minimize}}\quad\mathrm{MSE}\left[\widehat{B}\right]=H_{K}^{2}n_{b}^{4\gamma}+V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}},$ (23)
subject $\mbox{subject to}\quad-1/6<\gamma<0.$
Solving (23) gives $\gamma=-1/10$ and $3V_{K}\sigma^{2}(\theta_{0})=4H_{K}^{2}$. From Theorem 2, we can calculate $V_{K}\neq 0$ easily. Usually $\sigma^{2}(\theta_{0})>0$, so $H_{K}\neq 0$, which implies $D\neq 0$. This gives the rationality of $\alpha^{(5)}(\theta_{0})\neq 0$ in Assumption 3.
On the other hand, if $c_{1},...,c_{K}$ are randomly generated from $\mathcal{P}_{0}$, the optimization problem comes from the combination of (19) and (3.2.1):
$\underset{\gamma,\mathcal{P}_{0}}{\mbox{minimize}}\quad\mathbb{E}[H_{K}^{2}]n_{b}^{4\gamma}+\mathrm{Var}[H_{K}]n_{b}^{4\gamma}+\mathbb{E}[V_{K}]\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}},$ (24)
subject $\mbox{subject to}\quad-1/6<\gamma<0.$
Solving (24) also yields $\gamma=-1/10$. Thus, whether $c_{1},...,c_{K}$ are fixed or randomly generated, the optimal selection of $h_{k}$’s in the pilot stage is of order $O\left(n_{b}^{\gamma}\right)$ with $\gamma=-1/10$. It is noteworthy that the optimal perturbation in the pilot stage differs from the optimal one for the CFD problem shown in (8).
A more exact determination of $\{c_{1},...,c_{K}\}$ or $\mathcal{P}_{0}$ requires knowledge of $D$, which is more challenging than the estimation of $\alpha^{\prime}(\theta_{0})$. In this paper, we treat $\{c_{1},...,c_{K}\}$ and $\mathcal{P}_{0}$ as the hyperparameters whose selection is beyond the scope of our discussion.
4 Sample-recycling for Finite Difference Output
In Section 3, we have provided the estimators $\widehat{B}$ and $\widehat{\bm{\beta}}_{v}$ and shown that both of them are consistent. In this section, we use $\widehat{B}$ and $\widehat{\bm{\beta}}_{v}$ to set the estimated optimal perturbation by (8) and construct the estimators for $\alpha^{\prime}(\theta_{0})$.
The straightforward method is neglecting the $Kn_{b}$ pilot samples, i.e., setting
$\widehat{h}_{n_{2}}=\left(\frac{\widehat{\bm{\beta}}_{v}}{4n_{2}\widehat{B}^{2}}\right)^{1/6},$ (25)
where $n_{2}=n-Kn_{b}$. Then, $n_{2}$ sample pairs are used to construct $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ by (3) as the estimator of $\alpha^{\prime}(\theta_{0})$. From (9), given $\widehat{h}_{n_{2}}$, the expectation and variance of $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ are respectively
$\begin{cases}\mathbb{E}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\left|\widehat{h}_{n_{2}}\right.\right]=\alpha^{\prime}(\theta_{0})+B\left(\dfrac{\widehat{\bm{\beta}}_{v}}{4n_{2}\widehat{B}^{2}}\right)^{1/3}+o_{p}\left(n_{2}^{-1/3}\right),\\ \mathrm{Var}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\left|\widehat{h}_{n_{2}}\right.\right]=\bm{\beta}_{v}\left(\dfrac{\widehat{B}^{2}}{2n_{2}^{2}\widehat{\bm{\beta}}_{v}}\right)^{1/3}+o_{p}\left(n_{2}^{-2/3}\right).\end{cases}$ (26)
According to $\mathbb{E}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\right]=\mathbb{E}\left[\mathbb{E}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\left|\widehat{h}_{n_{2}}\right.\right]\right]$ and $\mathrm{Var}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\right]=\mathrm{Var}\left[\mathbb{E}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\left|\widehat{h}_{n_{2}}\right.\right]\right]+\mathbb{E}\left[\mathrm{Var}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\left|\widehat{h}_{n_{2}}\right.\right]\right]$, we have [^5] To be rigorous, (27) requires $o_{p}(\cdot)$ and $o_{p}^{2}(\cdot)$ in (26) to be uniformly integrable. Throughout the paper, we assume this condition implicitly when necessary to prevent excessive technical complexity. The rationality of the leading terms in (27) is referred to in Appendix B.1.
$\begin{cases}\mbox{Bias}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\right]=\left(\dfrac{B\bm{\beta}_{v}}{4n_{2}}\right)^{1/3}+o\left(n_{2}^{-1/3}\right),\\ \mathrm{Var}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\right]=\left(\dfrac{B^{2}\bm{\beta}_{v}^{2}}{2n_{2}^{2}}\right)^{1/3}+o\left(n_{2}^{-2/3}\right),\\ \mbox{MSE}\left[\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}\right]=3\left(\dfrac{B^{2}\bm{\beta}_{v}^{2}}{16n_{2}^{2}}\right)^{1/3}+o\left(n_{2}^{-2/3}\right).\end{cases}$ (27)
Li and Lam (2020) presents a similar result to (27). However, their approach to constructing $\widehat{h}_{n_{2}}$ differs from ours. For more details, interested readers may refer to Li and Lam (2020).
From (27), the MSE of $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ is $O\left(n_{2}^{-2/3}\right)$. When $n_{b}=o(n)$, the leading term of the MSE of $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ matches that of $\widehat{\Delta}_{n,h^{*}}$, where $h^{*}$ is the optimal perturbation as defined in (8). However, this result may not hold in some practical scenarios where the sample size is not large enough.
When $n_{b}\neq o(n)$ and $Kn_{b}=rn$, where $r$ is the proportion of the pilot samples and $10<r<$ 1, the MSE of $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ is
$3\left(\dfrac{B^{2}\bm{\beta}_{v}^{2}}{16(1-r)^{2}n^{2}}\right)^{1/3}+o\left(n^{-2/3}\right).$
Evidently, although this MSE is also of order $n^{2/3}$, it is larger than the optimal one in (9).
Therefore, the estimator $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ lacks efficiency because of its construction using only $n_{2}(<n)$ sample pairs. An intuitive way of improving the efficiency is to recycling the pilot samples in constructing the estimator. In the next section, we combine a pilot samples and the $n_{2}$ sample pairs to establish an efficient estimator.
4.1 Constructing the DSR-CFD estimator
In the following, we construct the double sample-recycling CFD (DSR-CFD) estimator through the remain $n_{2}$ sample pairs as well as the pilot sample pairs. We have to consider two questions: (1) how to generate the $n_{2}$ sample pairs; and (2) how to recycle the pilot sample pairs.
To answer the first question, note that we hope to utilize the whole $n$ sample pairs, including the pilot sample pairs and the $n_{2}$ sample pairs. Therefore, the desired CFD estimator should be constructed based on the perturbation $\widehat{h}_{n}$, where
$\widehat{h}_{n}=\left(\frac{\widehat{\bm{\beta}}_{v}}{4n\widehat{B}^{2}}\right)^{1/6}.$ (28)
Note that $\widehat{h}_{n}$ is derived from the entire sample-pair set of size $n$, whereas $\widehat{h}_{n_{2}}$, as defined in (25), is based on the subset of size $n_{2}$. Next, $n_{2}$ sample pairs are generated at $\widehat{h}_{n}$, and we can calculate $\left\{\Delta_{i}\left(\widehat{h}_{n}\right),i=1,...,n_{2}\right\}$ based on (2). When no confusion arises, the subscript $i$ of $\Delta_{i}\left(\widehat{h}_{n}\right)$ is omitted to denote a generic $\Delta\left(\widehat{h}_{n}\right)$.
To answer the second question, recall that our objective is to estimate $\alpha^{\prime}(\theta_{0})$ under the criteria of minimizing the MSE of its estimator. To reuse the pilot samples, we transform $\Delta(h_{k})$ which are constructed based on the pilot samples, by adjusting their location and scale, so that after the transformation, they have the same expectation and variance as $\Delta\left(\widehat{h}_{n}\right)$.
Specifically, for any $j=1,...,n_{b}$ and $k=1,...,K$, we transform $\Delta_{j}(h_{k})$ to
$\left(\dfrac{\mathrm{Var}\left[\Delta\left(\widehat{h}_{n}\right)\right]}{\mathrm{Var}[\Delta(h_{k})]}\right)^{1/2}[\Delta_{j}(h_{k})-\mathbb{E}\Delta(h_{k})]+\mathbb{E}\Delta\left(\widehat{h}_{n}\right),$ (29)
where $\mathrm{Var}\left[\Delta\left(\widehat{h}_{n}\right)\right]$ and $\mathrm{Var}[\Delta(h_{k})]$ are the variances at $\widehat{h}_{n}$ and $h_{k}$, respectively, and $\mathbb{E}\Delta\left(\widehat{h}_{n}\right)$ and $\mathbb{E}\Delta(h_{k})$ are the expectations at $\widehat{h}_{n}$ and $h_{k}$, respectively. Evidently, $[\Delta_{j}(h_{k})-\mathbb{E}\Delta(h_{k})]/\left(\mathrm{Var}[\Delta(h_{k})]\right)^{1/2}$ in (29) represents a standardization with mean 0 and variance 1. Then (29) possesses the same expectation and variance as $\Delta\left(\widehat{h}_{n}\right)$. Therefore, a desired estimator for $\alpha^{\prime}(\theta_{0})$ can be formulated as:
[…]
This is a weighted average of an estimator at $\widehat{h}_{n}$ constructed through the $n_{2}$ samples, and the one by the transformation from the pilot samples. It is reasonable to expect that under certain mild conditions, such as ensuring that the covariance of two terms in (30) converges faster than the variance of either term, the bias, variance and MSE of (30) all achieve their optimal values as in (9).
However, expectations and variances of $\Delta(h_{k})$ and $\Delta\left(\widehat{h}_{n}\right)$ in (29) are all unknown. Therefore, we have to substitute them in (30) by their estimators. In light of results in Section 3, we choose estimators as shown in Table 2.
Table 2: Estimators of expectations and variances of $\Delta(h_{k})$ and $\Delta\left(\widehat{h}_{n}\right)$.
Statistics Estimators
$\mathbb{E}\Delta(h_{k})$ $\left(1,h_{k}^{2}\right)\widehat{\beta}_{e}$
$\mathbb{E}\Delta\left(\widehat{h}_{n}\right)$ $\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\beta}_{e}$
$\mathrm{Var}[\Delta(h_{k})]$ ${\widehat{\bm{\beta}}_{v}^{2}}\big{/}{\left(2h_{k}^{2}\right)}$
$\mathrm{Var}\left[\Delta\left(\widehat{h}_{n}\right)\right]$ ${\widehat{\bm{\beta}}_{v}^{2}}\big{/}{\left(2{\widehat{h}_{n}}^{2}\right)}$
In Table 2, the estimators are all constructed through solutions to regression problems (14) and (16). Hence, they only depend on the pilot samples, and are independent with the remaining $n_{2}$ samples conditional on $\widehat{h}_{n}$, facilitating the theoretical analysis of the subsequently proposed estimator.
For any $j=1,...,n_{b}$ and $k=1,...,K$, we substitute these estimators into (29), and (29) is approximated by
$\Delta_{j}^{t}(h_{k})=\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\Delta_{j}(h_{k})-\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}\right)+\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}.$ (31)
Then, the DSR-CFD estimator, i.e., the estimated version of (30), is established as follows.
$\widehat{\Delta}_{n,\widehat{h}_{n}}=\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+\sum_{k=1}^{K}\sum_{j=1}^{n_{b}}\Delta_{j}^{t}(h_{k})\right).$ (32)
Next, we analyze the theoretical properties of $\widehat{\Delta}_{n,\widehat{h}_{n}}$. The results are summarized in the following theorem and proposition. To simplify the proof, we consider the non-random hyperparameters $c_{1},...,c_{K}$ such that when $n_{b}$ is given, $h_{1},...,h_{K}$ are also given. The proof is similar but more tedious when considering $\mathcal{P}_{0}$ as the hyperparameter, i.e., $c_{1},...,c_{K}$ are random.
Theorem 4.
Suppose that Assumptions 1, 2 and 3 hold. For any $k=1,...,K(K\geq 2)$, let $h_{k}=c_{k}n_{b}^{-1/10}(c_{k}\neq 0)$ and for any $j\neq k$, $c_{j}\neq c_{k}$. If $n_{b},n\to\infty$, then we have
$\mathbb{E}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=$ $\alpha^{\prime}(\theta_{0})+\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}n^{-1/3}+o\left(n^{-1/3}\right),$ (33)
$\mathrm{Var}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=$ $\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2n^{2}}\right)^{1/3}+\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2}\right)^{1/3}\frac{n_{b}}{n}\left[\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}-K\right]n^{-2/3}+o\left(n^{-2/3}\right),$ (34)
$\mathrm{MSE}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=$ $3\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/3}+\left\{\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/3}\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}\left(\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}+{\rm{sign}}(B)\sigma(\theta_{0})\right)\right.$
$+\left.\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2}\right)^{1/3}\frac{n_{b}}{n}\left[\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}-K\right]\right\}n^{-2/3}+o\left(n^{-2/3}\right),$ (35)
where ${\bm{c}}=[|c_{1}|,...,|c_{K}|]^{\top}$, ${\bm{c}^{4}}=[c_{1}^{4},...,c_{K}^{4}]^{\top}$, ${\bm{P}}=\bm{I}-\bm{X}_{e}(\bm{X}_{e}^{\top}\bm{X}_{e})^{-1}\bm{X}_{e}^{\top}$, $\bm{X}_{e}$ is defined as in (15), ${\rm{Diag}}(\bm{c^{-1}})={\rm{Diag}}\left(\dfrac{1}{|c_{1}|},...,\dfrac{1}{|c_{K}|}\right)$, ${\rm{sign}}(B)$ is the sign of $B$, and $||{\bm{v}}||_{2}=\left(\sum_{k=1}^{K}v_{k}^{2}\right)^{1/2}$ for any ${\bm{v}}\in\mathbb{R}^{K}$.
The proof of Theorem 4 is provided in Appendix B.2. Comparing (33), (34) and (35) with (9), the changes of bias, variance, and MSE in the DSR-CFD method compared to the optimal CFD method are evident, which depend on ${\bm{c^{\top}Pc^{4}}}$ and $\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}$. In Appendix B.3, we indicate that these changes are typically manageable.
Even more attractive is the fact that if $B$ and $D{\bm{c^{\top}Pc^{4}}}$ are of opposite signs, the squared bias of $\widehat{\Delta}_{n,\widehat{h}_{n}}$ will be smaller than that of the optimal CFD estimator (see (9)). In addition, when $c_{1},...,c_{K}$ are chosen such that $\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}\leq K$, the variance will be reduced. Proposition 5, a corollary of Theorem 4, further elaborates on these aspects.
Proposition 5.
Under the same conditions as in Theorem 4, we have
  • 1.
    If $n_{b}=o(n)$, then
    $\begin{cases}\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}=\alpha^{\prime}(\theta_{0})+\left(\dfrac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+o\left(n^{-1/3}\right),\\ \mathrm{Var}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{2n^{2}}\right)^{1/3}+o\left(n^{-2/3}\right),\\ \mathrm{MSE}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=3\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/3}+o\left(n^{-2/3}\right).\end{cases}$ (36)
  • 2.
    If $n_{b}=O(n)$, specifically, $\lim\limits_{n_{b},n\to\infty}Kn_{b}/n=r$, where $0<r\leq 1$, then
    • i.
      If $-1-{\rm{sign}}(B)\leq\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}D{\bm{c^{\top}Pc^{4}}}\leq 1-{\rm{sign}}(B)$, then
      $\left|\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}-\alpha^{\prime}(\theta_{0})\right|\leq\left(\dfrac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+o\left(n^{-1/3}\right).$
    • ii.
      Else, we have
      $\left|\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}-\alpha^{\prime}(\theta_{0})\right|\leq\left(1+\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}\left|D{\bm{c^{\top}Pc^{4}}}\right|\right)\left(\dfrac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+o\left(n^{-1/3}\right).$
  • 3.
    If $n_{b}=O(n)$, specifically, $\lim\limits_{n_{b},n\to\infty}Kn_{b}/n=r$, where $0<r\leq 1$, and $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$, then
    $\mathrm{Var}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]\leq\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{2n^{2}}\right)^{1/3}+o\left(n^{-2/3}\right).$
These results in Proposition 5 can be easily derived by the same technique as in Theorem 4 and the proof is provided in Appendix B.4. Proposition 5 emphasizes that the DSR-CFD estimator possesses a reduced variance and in some cases, a reduced bias, compared to the optimal CFD estimator, and it is applicable even if $r=1$. Specifically, Proposition 5 guides the following conclusions:
  1. 1.
    The first result indicates that the DSR-CFD method regains the CFD method when $n_{b}=o(n)$ (refer to (9) and (36)). This aligns with intuition since we allocate a majority of the samples to gradient estimation rather than constant estimation.
  2. 2.
    The second result in Proposition 5 suggests that if $n_{b}=O(n)$, then the absolute value of the bias of the DSR-CFD estimator converges at the same rate as that of the optimal CFD estimator, or it may even be reduced compared to the latter. In particular, 2.i in Proposition 5 demonstrates that the former is smaller than the latter. On the other hand, in Appendix B.3, we show that ${\bm{c^{\top}Pc^{4}}}$ is typically controllable. Therefore, 2.ii in Proposition 5 illustrates that even if the absolute value of the bias of the DSR-CFD estimator is not reduced, its increment is moderate and will not blow up.
  3. 3.
    The third result in Proposition 5 also demonstrates that if $n_{b}=O(n)$ and $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$, the variance of the DSR-CFD estimator is reduced compared to that of the optimal CFD estimator. It is worth mentioning that the condition $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$ is a highly sufficient condition of $\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}\leq K$, and the latter leads to a reduced variance (see Theorem 4).
  4. 4.
    An appealing observation from Proposition 5 is the robustness of the DSR-CFD method w.r.t. $r$, in the sense that our proposed DSR-CFD estimator exhibits favorable theoretical performance for both $r=0$ and $r>0$. In particular, when $r=1$, i.e., indicating that the entire simulation budget is allocated to the pilot phase, the DSR-CFD estimator remains efficient. This theoretical insight serves as a guide, suggesting that when faced with limited simulation resources, allocating the entire budget to generate pilot samples, and using the double sample-recycling technique can lead to an efficient finite difference estimator.
Remark 2.
The efficiency of the DSR-CFD estimator is attributed to the expectation and variance of (29) being identical to those of the optimal CFD estimator. However, by substituting $\mathbb{E}[\Delta(h_{k})]$, $\mathbb{E}\left[\Delta\left(\widehat{h}_{n}\right)\right]$, $\mathrm{Var}[\Delta(h_{k})]$ and $\mathrm{Var}\left[\Delta\left(\widehat{h}_{n}\right)\right]$ using their regression estimators, respectively, so we introduce the intrinsic correlation and this may reduce the variance as well as the bias. In fact, the reason for variance reduction and possible bias reduction lies in the substitution of $\mathbb{E}[\Delta(h_{k})],k=1,...,K$. If we impractically assume that $\mathbb{E}\Delta(h_{k}),k=1,...,K$ are all known precisely and denote $\Delta_{j}^{t}(h_{k})$ by
$\Delta_{j}^{t}(h_{k})=\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\Delta_{j}(h_{k})-\mathbb{E}\Delta(h_{k})\right)+\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e},$
then the new DSR-CFD estimator, denoted by $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$, fully recovers the result of the optimal CFD estimator. In Appendix B.5, we provide these properties. After substituting $\mathbb{E}[\Delta(h_{k})],k=1,...,K$ with their regression estimators $\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e},k=1,...,K$, two scenarios may occur:
  • Variance reduction: For any $k=1,...,K$, $\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}$ is highly positively correlated with $\Delta_{j}(h_{k})$ since $\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}$ is an approximation of $\Delta_{j}(h_{k})$. Although $\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}$ is not zero-mean, it can be seen as a control variate because only the variance is considered. In the end, these control variates will reduce the variance (see Glasserman, 2013).
  • Possible bias reduction: As observed in (15), the regression problem is non-canonical and $\mathbb{E}[\mathcal{E}_{e}]\neq 0$. Therefore, $\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}$ is a biased estimator of $\mathbb{E}[\Delta(h_{k})]$ which leads to the difference between the biases. It is more obvious if we rewrite the DSR-CFD estimator as
    $\widehat{\Delta}_{n,\widehat{h}_{n}}=$ $\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+n_{b}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\widehat{\Delta}_{n_{b},h_{k}}-\mathbb{E}\Delta(h_{k})\right)+Kn_{b}\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}\right)$
    $+\frac{n_{b}}{n}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\mathbb{E}\Delta(h_{k})-\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}\right)$
    $=$ $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}+\frac{n_{b}}{n}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\mathbb{E}\Delta(h_{k})-\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}\right).$
    Note that this bias adjustment depends on $D$ (see (15)) while the bias of the optimal CFD estimator depends on $B$ (see (9)). If these two bias terms have a tendency to cancel each other out, the increment of $\mathrm{Bias}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]$ may pull $\mathrm{Bias}\left[\widehat{\Delta}_{n,h^{*}}\right]$ towards 0, thus reducing the bias.
In Section 6, we will present several numerical examples to validate the aforementioned properties of our proposed DSR-CFD estimator.
5 Implementation
Previous sections have provided efficient estimators for $\bm{\beta}_{e}$ and $\bm{\beta}_{v}$ and constructed a double sample-recycling based estimator. The theoretical bias, variance and MSE have also been calculated, that is Theorem 4. This section will provide the details for estimating $\bm{\beta}_{e}$ and $\bm{\beta}_{v}$ and constructing the DSR-CFD estimator.
5.1 Estimating parameters $B$ and $\sigma^{2}(\theta_{0})$
In this subsection, we present the specific steps for estimating unknown constants using the bootstrap method as described in Section 3. We generate i.i.d. $c_{1},...,c_{K}$ from $\mathcal{P}_{0}$ and for $k=1,...,K$, let $h_{k}=c_{k}n_{b}^{-1/10}$. At each perturbation $h_{k}$, we generate $n_{b}$ samples $\bm{\Delta}(h_{k})=\{\Delta_{1}(h_{k}),...,\Delta_{n_{b}}(h_{k})\}$. Repeat independently the bootstrap procedure in Section 3 for $I$ times, we obtain $I$ versions of $\widehat{\Delta}^{b}_{n_{b},h_{k}}$, denoted by $\left\{\widehat{\Delta}^{b}_{n_{b},h_{k}}(1),...,\widehat{\Delta}^{b}_{n_{b},h_{k}}(I)\right\}$. Then $\mathbb{E}_{*}\widehat{\Delta}_{n_{b},h_{k}}$ is approximated by
$\bar{\Delta}^{b}_{n_{b},h_{k}}=\frac{1}{I}\sum_{q=1}^{I}\widehat{\Delta}^{b}_{n_{b},h_{k}}(q),$ (37)
and $\mathrm{Var}_{*}\left[\widehat{\Delta}_{n_{b},h_{k}}\right]$ is approximated by
$s^{2}_{n_{b},h_{k}}=\frac{1}{I}\sum_{q=1}^{I}\left[\widehat{\Delta}^{b}_{n_{b},h_{k}}(q)-\bar{\Delta}^{b}_{n_{b},h_{k}}\right]^{2}.$ (38)
According to the law of large numbers, as $I$ increases, $\bar{\Delta}^{b}_{n_{b},h_{k}}$ and $s^{2}_{n_{b},h_{k}}$ are respectively consistent estimators of $\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h_{k}}$ and $\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h_{k}}\right]$.
It is worth noting that both of the regression problems (14) and (16) exhibit heteroscedasticity. This arises from the fact that elements of $\mathcal{E}_{e}$ in (15) and $\mathcal{E}_{v}$ in (17) possess varying variances, due to different values of $h_{k}$’s. The primary approach to addressing heteroscedasticity is through the weighted least squares method. This involves determining appropriate weights to eliminate or mitigate the heteroscedasticity in the regression models, i.e., (14) and (16). As the estimator obtained through the weighted least squares method exhibits the same convergence rate as that derived from the ordinary least squares method (Hayashi, 2011), we omit an in-depth analysis of the weighted least squares method.
While estimating the parameter $B$, we can effectively tackle heteroscedasticity by estimating the variance at each perturbation using the $n_{b}$ samples, and the reciprocals of the standard deviations are appropriate weights. That is, we regress $\widetilde{\bm{Y}}_{e}$ on $\widetilde{\bm{X}}_{e}$, where
$\widetilde{\bm{Y}}_{e}$ $=\left[\bar{\Delta}^{b}_{n_{b},h_{1}}\big{/}s_{n_{b},h_{1}},...,\bar{\Delta}^{b}_{n_{b},h_{K}}\big{/}s_{n_{b},h_{K}}\right]^{\top},\widetilde{\bm{X}}_{e}=\left[\begin{array}[]{lll}1\big{/}s_{n_{b},h_{1}}&...&1\big{/}s_{n_{b},h_{K}}\\ h_{1}^{2}\big{/}s_{n_{b},h_{1}}&...&h_{K}^{2}\big{/}s_{n_{b},h_{K}}\end{array}\right]^{\top},$ (41)
and the slope is the estimator $\widehat{B}$.
However, when estimating $\sigma^{2}(\theta_{0})$, we lack such standard deviations. From $\mathcal{E}_{e}$ in (17), it can be seen that the heteroscedasticity of the regression model (16) arises from the different variances of $\phi(\cdot)$ for different values of $h_{k}$’s. It follows from (13) that $h_{k}^{2}\phi(h_{k})$ ($k=1,...,K$) share the same variance. Therefore, we can handle the heteroscedasticity resulting from differences in perturbations by multiplying both sides of (16) by $h_{k}^{2}$ ($k=1,...,K$) simultaneously. That is, we regress $\widetilde{\bm{Y}}_{v}$ on $\widetilde{\bm{X}}_{v}$, where
$\widetilde{\bm{Y}}_{v}=\left[h_{1}^{2}s^{2}_{n_{b},h_{1}},...,h_{K}^{2}s^{2}_{n_{b},h_{K}}\right]^{\top},\widetilde{\bm{X}}_{v}=\left[\frac{n_{b}-1}{2n_{b}^{2}},...,\frac{n_{b}-1}{2n_{b}^{2}}\right]^{\top},$ (42)
and the slope is the estimator $\widehat{\bm{\beta}}_{v}$.
5.2 Estimating $\alpha^{\prime}(\theta_{0})$
Once we have generated $\Delta_{j}(h_{k})$, where $j=1,2,...,n_{b}$ and $k=1,2,...,K$, and have estimated the unknown constants using these samples, we can transform these samples using (31) and construct the DSR-CFD estimator using (32). We summarize the above procedure as Algorithm 1:
  1. 1.
    Usually, we choose $\mathcal{P}_{0}$ as a truncated normal distribution $\mathcal{P}_{0}=\psi\left(\mu_{0},\sigma_{0}^{2},L,U\right)$, where $0<L<U<\infty$. This setting ensures that $\mathbb{E}[H_{K}]$, $\mathrm{Var}[H_{K}]$, $\mathbb{E}[V_{K}]$, $\mathbb{E}\left[\widetilde{H}_{K}\right]$, $\mathrm{Var}\left[\widetilde{H}_{K}\right]$, $\mathbb{E}\left[\widetilde{V}_{K}\right]$, $\mathbb{E}\left[\widehat{H}_{K}\right]$,$\mathrm{Var}\left[\widehat{H}_{K}\right]$ and $\mathbb{E}\left[\widehat{V}_{K}\right]$ are all bounded. Note that for $c_{1},...,c_{K}\sim\mathcal{P}_{0}$, the condition $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$ holds only when $U\leq\sqrt{2}L$, which is a stringent condition. However, we have mentioned that $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$ is a sufficient condition of $\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}\leq K$ which is satisfied more easily for $\mathcal{P}_{0}$ with different parameters.
  2. 2.
    In step 1, the pilot samples are generated at the perturbations $h_{k}=c_{k}n_{b}^{-1/10}$, rather than $c_{k}n_{b}^{-1/6}$. This is because (23) and (24) indicate that in the pilot stage, the optimal perturbation is of order $n_{b}^{-1/10}$.
  3. 3.
    In step 1, when estimating $B$ and $\sigma^{2}(\theta_{0})$, we apply the weighted least squares regression proposed in Section 5.1.
  4. 4.
    In step 2, we generate the remaining samples at $\widehat{h}_{n}$ defined as in (28), instead of $\widehat{h}_{n_{2}}$ defined as in (25), because we will utilize the total $n$ samples to construct the CFD estimator.
Input: The number of total sample pairs $n$, the number of perturbation parameters $K$, the number of pilot sample pairs at each perturbation $n_{b}$, the number of bootstrap $I$. Initial perturbation generator $\mathcal{P}_{0}$.
Step 1. Estimate the unknown constants $B$, $\sigma^{2}(\theta_{0})$ and the optimal perturbation.
  1. 1.
    Generate $c_{1},c_{2},...,c_{K}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{P}_{0}$ and let $h_{k}=c_{k}n_{b}^{-1/10}$ for any $k=1,2,...,K$. Generate pilot samples and calculate $\Delta_{j}(h_{k})$, where $j=1,2,...,n_{b}$ and $k=1,2,...,K$.
  2. 2.
    Perform bootstrap resampling $I$ times for each set of samples at $h_{1},h_{2},...,h_{K}$ and calculate $\bar{\Delta}^{b}_{n_{b},h_{k}}$ and $s^{2}_{n_{b},h_{k}}$ for $k=1,...,K$, according to (37) and (38).
  3. 3.
    Regress $\widetilde{\bm{Y}}_{e}$ on $\widetilde{\bm{X}}_{e}$, where $\widetilde{\bm{Y}}_{e}$ and $\widetilde{\bm{X}}_{e}$ are defined as in (41), and the slope is $\widehat{B}$.
    Regress $\widetilde{\bm{Y}}_{v}$ on $\widetilde{\bm{X}}_{v}$, where $\widetilde{\bm{Y}}_{v}$ and $\widetilde{\bm{X}}_{v}$ are defined as in (42), and the slope is $\widehat{\bm{\beta}}_{v}$.
    Set the perturbation by (28), i.e.,
    $\widehat{h}_{n}=\left(\frac{\widehat{\bm{\beta}}_{v}}{4n\widehat{B}^{2}}\right)^{1/6}.$
Step 2. Construct the DSR-CFD estimator of $\alpha^{\prime}(\theta_{0})$.
  1. 1.
    Reuse $\Delta_{j}(h_{k})$, where $j=1,2,...,n_{b}$ and $k=1,2,...,K$. That is, calculate $\Delta_{j}^{t}(h_{k})=\dfrac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\Delta_{j}(h_{k})-\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}\right)+\left(1,{\widehat{h}}_{n}^{2}\right)\widehat{\bm{\beta}}_{e}$ for $j=1,2,...,n_{b}$, and $k=1,2,...,K$.
  2. 2.
    Denote the number of remaining sample pairs as $n_{2}=n-Kn_{b}$. Generate $n_{2}$ sample pairs at $\widehat{h}_{n}$ and calculate $\left\{\Delta_{1}(\widehat{h}_{n}),...,\Delta_{n_{2}}(\widehat{h}_{n})\right\}$.
  3. 3.
    The DSR-CFD estimator
    $\widehat{\Delta}_{n,\widehat{h}_{n}}=\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+\sum_{k=1}^{K}\sum_{j=1}^{n_{b}}\Delta_{j}^{t}(h_{k})\right).$
Output: $\widehat{\Delta}_{n,\widehat{h}_{n}}$.
Algorithm 1 (DSR-CFD method.).
6 Numerical Experiments
In this section, we use three examples to check the performance of our method. Examples 1 and 2 are from Li and Lam (2020). Example 3 is the generic $M/M/1$ queueing system from Lam et al. (2022). All numerical results are based on 1000 replications.
We will compare the performance of our proposed method and the EM-CFD method proposed by Li and Lam (2020). The detailed settings of our proposed method are presented as follows:
  • In the first and third examples, we let the initial perturbation generator $\mathcal{P}_{0}=\psi\left(0,1,0.1,\infty\right)$ ($\psi\left(0,0.1,0.01,\infty\right)$ in Example 2), the number of bootstrap $I=1000$. During the first phase, we set the number of regression points as $K=10$ or $20$.
  • The bootstrap-based method, labeled by “BOOT-CFD”, represents using the bootstrap method to estimate the unknown constants but not reusing the pilot samples.
Example 1.
Consider estimating the first order derivative of the function $\alpha(\theta)=k{\rm sin}(\theta)$ at the point $\theta_{0}=0$, where $k$ is a priori unknown parameter.
  • Case 1: The observed variables at $\theta$ are assumed to obey a normal distribution with mean $\alpha(\theta)$ and variance $1$, i.e., $Y(\theta)\sim\mathcal{N}(\alpha(\theta),1)$.
  • Case 2: The observed variables at $\theta$ are assumed to obey a normal distribution with mean $\alpha(\theta)$ and variance ${\rm e}^{-3\theta}$, i.e., $Y(\theta)\sim\mathcal{N}(\alpha(\theta),{\rm e}^{-3\theta})$.
Cases 1 and 2 represent homoscedasticity and heteroscedasticity, respectively. From a simple calculation we obtain
$\alpha^{\prime}(\theta_{0})=k,\quad B=-k/6,\quad\sigma^{2}(\theta_{0})=1,$
as the real parameters to measure the performance of our proposed method. During the execution of each algorithm, we assume that the parameters $\alpha^{\prime}(\theta_{0})$, $B$ and $\sigma^{2}(\theta_{0})$ are unknown.
Initially, we assess the validity and robustness of the bootstrap method for estimating unknown constants and perturbations. In the implementation, we set $k=10$, $K=10$ and utilize 200 pilot sample pairs. The results, presented in Figures 1 and 2, display three subfigures from left to right, representing the estimators of $B$, $\sigma^{2}(\theta_{0})$, and $h$ across 1000 replications. When executing the EM-CFD algorithm, to align with the setting in Li and Lam (2020), we generate $h_{1},\ldots,h_{200}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,1\times 200^{-1/5}\right)$.
In the left subfigure of Figure 1, we observe that our method’s robustness in capturing third-order derivative information is slightly lower than that of the EM-CFD algorithm in case 1 since DSR-CFD method has a wider box. This happens because the DSR-CFD method has only $K=10$ regression points instead of 200. Nonetheless, the DSR-CFD method is still comparable due to the reduced error at each regression point. The middle subfigure of Figure 1 indicates our approach effectively estimates variance information, a challenging task for the EM-CFD algorithm as it incorporates the variance of the perturbation generator. The right subfigure provides a comprehensive view of perturbation estimation. Across all 1000 experiments, our method consistently outperforms the EM-CFD algorithm, approaching the true value more closely.
Figure 1: Comparison of the estimates of the unknown constants and perturbation for case 1 in Example 1 $(k=10)$.
In Figure 2, we illustrate the impact of various approaches on the estimation of $B$, $\sigma^{2}(\theta_{0})$, and $h$ in case 2 of Example 1, which deals with scenarios of heteroskedasticity. Analyzing the pair of subfigures presented on the left, it is apparent that our method captures the variance information effectively and provides a more robust estimate of $B$. From the right subfigure, we observe that our method yields a more accurate and reliable estimator of the perturbation. In contrast, the perturbations obtained by the EM-CFD algorithm deviate significantly from the true value, exhibit higher fluctuations.
Figure 2: Comparison of the estimates of the unknown constants and perturbation for case 2 in Example 1 $(k=10)$.
In Figure 3, we depict the sensitivity of different methods to $r$, where the suffixes “1” and “2” denote cases 1 and 2, respectively, in Example 1. We set the number of sample pairs to 1000. When examining the EM-CFD method, we generate $h_{1},\ldots,h_{1000r}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,1\times(1000r)^{-1/5}\right)$. As observed in the figure, the presence of heteroskedasticity complicates the problem for each method, resulting in inferior estimations. It is worth mentioning that the impact of heteroskedasticity is more pronounced in EM-CFD. When $r<0.55$, the EM-CFD method exhibits significant instability. The DSR-CFD method consistently outperforms other methods in various scenarios, surpassing the EM-CFD method in both simple and complex situations. Additionally, the DSR-CFD method demonstrates stability concerning $r$, while the performance of EM-CFD and BOOT-CFD worsens with increasing $r$. Even when $r=1$, the DSR-CFD method maintains strong performance, making it suitable for small sample sizes, where all samples are utilized for both constant and gradient estimation, effectively leveraging the sample information.
Figure 3: Illustration of the MSE of different methods with $r$ for cases 1 and 2 in Example 1 $(k=10)$.
Example 2.
We consider estimating the first order derivatives of the polynomial function $\alpha(\theta)=1-6\theta+36\theta^{2}-53\theta^{3}+22\theta^{5}$ at different $\theta_{0}$. Assume that the observed variables at $\theta$ obey a normal distribution with mean $\alpha(\theta)$ and variance $0.05$, i.e., $Y(\theta)\sim\mathcal{N}(\alpha(\theta),0.05)$. From a simple calculation we obtain
$\alpha^{\prime}(\theta_{0})=-6+72\theta_{0}-159\theta_{0}^{2}+110\theta_{0}^{4},\quad B=-53+220\theta_{0}^{2},\quad\sigma^{2}(\theta_{0})=0.05,$
as the real parameters to measure the performance of our proposed method. During the implementation of each algorithm, we assume that the parameters $\alpha^{\prime}(\theta_{0})$, $B$ and $\sigma^{2}(\theta_{0})$ are unknown.
Table 3 presents a comparison of bias, variance, and MSE between the DSR-CFD estimator (denoted by Bias-DSR, Var-DSR and MSE-DSR, respectively) and the Opt-CFD estimator (denoted by Bias-Opt, Var-Opt and MSE-Opt, respectively) across various sample sizes and estimation points (corresponding to different problems) in example 2. In Table 3, $B$ and $D$ are respectively the true values and Opt-CFD is meant to set $B$ and $\sigma^{2}(\theta_{0})$ as their true values, i.e., $B=-53+220\theta_{0}^{2}$ and $\sigma^{2}(\theta_{0})=0.05$. We take 50% of the total sample to estimate the unknown constants and set $K=10$ for DSR-CFD method.
Table 3: Comparison of the bias, variance and MSE of the DSR-CFD and Opt-CFD estimators in Example 2.
Budget $x$ $B$ $D$ Bias-DSR Bias-Opt Var-DSR Var-Opt MSE-DSR MSE-Opt
$10^{3}$ 0 $-53$ 22 $-0.1335$ $-0.0829$ 0.0117 0.0154 0.0295 0.0222
0.2 $-44.2$ 22 $-0.1356$ $-0.0789$ 0.0115 0.0149 0.0299 0.0211
0.4 $-17.5$ 22 $-0.1264$ $-0.0584$ 0.0105 0.0067 0.0265 0.0101
0.6 26.2 22 0.0170 0.0693 0.0102 0.0101 0.0105 0.0149
0.8 87.8 22 0.0658 0.1016 0.0163 0.0212 0.0206 0.0315
1 167 22 0.0991 0.1294 0.0217 0.0323 0.0315 0.0490
$10^{4}$ 0 $-53$ 22 $-0.0583$ $-0.0415$ 0.0021 0.0031 0.0055 0.0049
0.2 $-44.2$ 22 $-0.0558$ $-0.0339$ 0.0021 0.0031 0.0052 0.0042
0.4 $-17.5$ 22 $-0.0544$ $-0.0279$ 0.0017 0.0015 0.0047 0.0023
0.6 26.2 22 0.0099 0.0343 0.0018 0.0021 0.0019 0.0033
0.8 87.8 22 0.0340 0.0484 0.0033 0.0044 0.0044 0.0068
1 167 22 0.0549 0.0596 0.0055 0.0069 0.0085 0.0104
$10^{5}$ 0 $-53$ 22 $-0.0245$ $-0.0196$ 4.5424e-04 6.7200e-04 0.0011 0.0011
0.2 $-44.2$ 22 $-0.0222$ $-0.0179$ 4.2904e-04 6.3269e-04 9.2243e-04 9.5076e-04
0.4 $-17.5$ 22 $-0.0208$ $-0.0126$ 2.8572e-04 3.3534e-04 7.1882e-04 4.9412e-04
0.6 26.2 22 0.0088 0.0150 3.4222e-04 4.2576e-04 4.1963e-04 6.4935e-04
0.8 87.8 22 0.0200 0.0220 7.4610e-04 9.9544e-04 0.0011 0.0015
1 167 22 0.0270 0.0277 9.8655e-04 0.0015 0.0017 0.0023
From Table 3, we have the following observations:
  • Bias: The bias is influenced by the signs of $B$ and $D$. If $B$ and $D$ have the same (opposite) signs, the absolute bias of the DSR-CFD method is smaller (larger) than that of the Opt-CFD method. This aligns with Theorem 4 because $\bm{c^{\top}Pc^{4}}$ is negative. When $B$ and $D$ share the same (opposite) signs, $B$ and $D\bm{c^{\top}Pc^{4}}$ exhibit opposite (same) signs. From Proposition 5, the ratio of the abstract values of biases of the DSR-CFD estimator and the Opt-CFD estimator is bounded by 1 or $1+\sqrt{\frac{r}{K}}\frac{2}{\sigma(\theta_{0})}\left|D{\bm{c^{\top}Pc^{4}}}\right|$.
[…]
  • Variance: It is observed that as the sample size increases, the variance of the DSR-CFD estimator consistently outperforms that of the Opt-CFD estimator, aligning with Proposition 5. For instance, at $x=1$ with a sample budget of $10^{5}$, the variances for the DSR-CFD and Opt-CFD estimators are $9.8655\times 10^{-4}$ and $0.0015$, respectively.
  • MSE: It is important to note that the bias and variance of the DSR-CFD estimator consistently approximate and are smaller than those of the Opt-CFD estimator, respectively. Consequently, when employing MSE as the error criterion, the DSR-CFD estimator consistently performs nearly as well as (or even outperforms) the Opt-CFD estimator.
We examine the estimated MSE in Example 2 by setting $\theta_{0}=0$ and varying the sample pairs from $n=50$ to $n=500$. When considering the EM-CFD algorithm, we set $r=0.1$ and generate $h_{1},...,h_{rn}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,0.01\times(rn)^{-1/5}\right)$. When applying the DSR-CFD algorithm, we set $K=10$ and $r=1$. That is, the entire simulation budget is allocated to the pilot phase. Figure 4 shows the results. From the figure, we observe that the estimated MSE of the EM-CFD estimator decreases gradually as the sample size increases. Compared to the Opt-CFD and DSR-CFD estimators, however, the EM-CFD estimator exhibits a significantly larger MSE. Moreover, the MSE of the DSR-CFD estimator is consistently close to that of the Opt-CFD estimator. For example, when the sample size is 50, the MSE for the EM-CFD method is approximately 5.959, while the MSEs for the DSR-CFD and Opt-CFD methods are approximately 0.421 and 0.170, respectively.
Figure 4: Variation of the estimated MSE by different methods with increasing sample size in Example 2 $(\theta_{0}=0)$.
Example 3.
We consider a generic $M/M/1$ queueing system which is empty initially. Both of the arrival distribution and the service distribution are the exponential distribution and the arrival rate and service rate are denoted by $\lambda$ and $\mu$, respectively. Observations are the averaged system time of the first $N$ customers and we are interested in the gradient of this quantity w.r.t. the service rate. For this example, we consider two cases.
  • Case 1: We consider a critically loaded system, where we set $\lambda=\mu=4$ and $N=10$. In this case, the true derivative is $-0.2501$, which is calculated using the likelihood ratio/score function method (see e.g. Glynn, 1990) with $10^{6}$ simulation repetitions (Lam et al., 2022).
  • Case 2: We consider a non-critically loaded queue, where we set $\lambda=3$, $\mu=5$ and $N=10$. In this case, the true derivative is $-0.1136$.
In this example, we denote the estimate of $B$ as $\widetilde{B}$ in Appendix B.1 and set $\epsilon=5\times 10^{-4}$. Figure 5 illustrates the numerical results for different methods in Example 3 when the sample-pair size changes from $n=60$ to $n=1000$. We compare the following three methods: The conventional central finite difference (Con-CFD) method, EM-CFD method and DSR-CFD method. For the Con-CFD method, we do not know any information about the model in advance and set $B=5$ and $\sigma^{2}(\theta_{0})=1$. When considering the EM-CFD method, we set $r=0.1$ and generate $h_{1},...,h_{rn}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,1\times(rn)^{-1/5}\right)$. When applying the DSR-CFD method, we set $K=20$ in the pilot phase and $r=1$.
Figure 5: Comparison of the MSE among different estimators for cases 1 and 2 in Example 3.
Figure 5 illustrates that the performance of the Con-CFD estimator is inferior compared to the other two estimators. This observation emphasizes the importance of estimating unknown constants within finite difference methods. In case 1 of Example 3, the DSR-CFD estimator consistently yields the most favorable results. For instance, with a sample-pair size of 60, the MSEs for the DSR-CFD and EM-CFD estimators are $0.011$ and $0.033$, respectively. In case 2, the performances of these two estimators are similar. In conclusion, the DSR-CFD estimator performs effectively across various problems.
7 Conclusions
The choice of the perturbation significantly impacts the accuracy of the finite difference approximation in stochastic gradient estimation. Although the order of the optimal perturbation is known, determining the unknown constant remains a challenge. To address this issue and construct an efficient finite difference approximation, in this paper, we have developed a double sample-recycling approach aimed at maximizing sample efficiency. Theoretical analyses reveal that the proposed estimator exhibits a reduced variance compared to the optimal finite difference estimator, and in some cases, a decrease in bias. In particular, we provide algorithms and numerical experiments. The theory and the findings of the numerical experiments are consistent.
Appendix A A Proofs in Section 3
A.1 Proof of Theorem 1
We prove (11) and (12) separately.
[…]
$\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h}=\mathbb{E}_{*}[\Delta_{1}^{*}(h)]=\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}\Delta_{i}(h)=\widetilde{\alpha}_{h}^{\prime}(\theta_{0})+\bar{Z}_{n_{b},h},$ (43)
where the first equality is due to the linearity of expectation, the second one is due to Table 1 and the last one is from (4). Under Assumption 3, using the Taylor expansion, we get
$\alpha(\theta_{0}+h)=\alpha(\theta_{0})+\alpha^{\prime}(\theta_{0})h+\frac{\alpha^{(2)}(\theta_{0})}{2}{h}^{2}+B{h}^{3}+\frac{\alpha^{(4)}(\theta_{0})}{24}h^{4}+\frac{\alpha^{(5)}(\theta_{0})}{120}h^{5}+o(h^{5}),$
$\alpha(\theta_{0}-h)=\alpha(\theta_{0})-\alpha^{\prime}(\theta_{0})h+\frac{\alpha^{(2)}(\theta_{0})}{2}h^{2}-Bh^{3}+\frac{\alpha^{(4)}(\theta_{0})}{24}h^{4}-\frac{\alpha^{(5)}(\theta_{0})}{120}h^{5}+o(h^{5}).$
Subtraction eliminates $\alpha^{(2)}(\theta_{0})$ and $\alpha^{(4)}(\theta_{0})$, leaving
$\widetilde{\alpha}_{h}^{\prime}(\theta_{0})=\alpha^{\prime}(\theta_{0})+Bh^{2}+Dh^{4}+o(h^{4}),$
where $D=\alpha^{(5)}(\theta_{0})/120$. Combining with (43) gives
$\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h}=\alpha^{\prime}(\theta_{0})+Bh^{2}+Dh^{4}+o(h^{4})+\bar{Z}_{n_{b},h}.$
[…]
Under $\mathbb{P}_{*}$, $\Delta_{1}^{*}(h),\Delta_{2}^{*}(h),...,\Delta_{n_{b}}^{*}(h)$ are i.i.d., so the variance of $\widehat{\Delta}^{b}_{n_{b},h}$ is
$\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]$ $=\frac{1}{n_{b}}\mathrm{Var}_{*}[\Delta_{1}^{*}(h)]=\frac{1}{n_{b}}\left(\mathbb{E}_{*}\left[\Delta_{1}^{*}(h)\right]^{2}-\left[\mathbb{E}_{*}\Delta_{1}^{*}(h)\right]^{2}\right)$
$=\frac{1}{n_{b}}\left(\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}\Delta_{i}^{2}(h)-\left(\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}\Delta_{i}(h)\right)^{2}\right)$
$=\frac{1}{n_{b}^{2}}\sum_{i=1}^{n_{b}}\left(\Delta_{i}(h)-\frac{1}{n_{b}}\sum_{j=1}^{n_{b}}\Delta_{j}(h)\right)^{2}.$
Denote
$S^{2}=\frac{1}{n_{b}-1}\sum_{i=1}^{n_{b}}\left(\Delta_{i}(h)-\frac{1}{n_{b}}\sum_{j=1}^{n_{b}}\Delta_{j}(h)\right)^{2}=\frac{1}{n_{b}-1}\sum_{i=1}^{n_{b}}\left(Z_{i}(h)-\bar{Z}_{n_{b},h}\right)^{2},$ (44)
and then we have
$\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]=\frac{n_{b}-1}{n_{b}^{2}}S^{2}.$ (45)
In the following we study $S^{2}$. Since $Z_{1}(h),Z_{2}(h),...,Z_{n_{b}}(h)$ are i.i.d., $S^{2}$ is an unbiased estimate of $\mathrm{Var}[Z_{1}(h)]$, i.e.,
$\mathbb{E}[S^{2}]=\mathrm{Var}[Z_{1}(h)]=\mathrm{Var}[\Delta_{1}(h)]=\frac{\sigma^{2}(\theta_{0}+h)+\sigma^{2}(\theta_{0}-h)}{4h^{2}}.$
Under Assumption 2, we have $\sigma^{2}(\theta_{0}+h)=\sigma^{2}(\theta_{0})+o(1)$ and $\sigma^{2}(\theta_{0}-h)=\sigma^{2}(\theta_{0})+o(1)$. Therefore,
$\mathbb{E}[S^{2}]=\mathrm{Var}[Z_{1}(h)]=\frac{\sigma^{2}(\theta_{0})+o(1)}{2h^{2}}.$ (46)
Let $\xi(h)$ is the zero-mean error associated with $S^{2}$ and we have
$S^{2}=\frac{\sigma^{2}(\theta_{0})+o(1)}{2h^{2}}+\xi(h).$ (47)
The variance of $\xi(h)$ is
$\mathrm{Var}[\xi(h)]=\mathbb{E}[\xi(h)]^{2}=\mathbb{E}\left[S^{2}-\frac{\sigma^{2}(\theta_{0})+o(1)}{2h^{2}}\right]^{2}=\mathbb{E}[S^{2}]^{2}-\frac{\sigma^{4}(\theta_{0})+o(1)}{4h^{4}}.$ (48)
We calculate $\mathbb{E}[S^{2}]^{2}$ as follows. It follows from (44) that
$\quad\ (n_{b}-1)^{2}\mathbb{E}[S^{2}]^{2}$
$=\mathbb{E}\left[\sum_{i=1}^{n_{b}}\left(Z_{i}(h)-\bar{Z}_{n_{b},h}\right)^{2}\right]^{2}=\mathbb{E}\left[\sum_{i=1}^{n_{b}}Z_{i}^{2}(h)-n_{b}\bar{Z}_{n_{b}}^{2}(h)\right]^{2}$
$=\mathbb{E}\left[\sum_{i=1}^{n_{b}}Z_{i}^{2}(h)\right]^{2}-2n_{b}\mathbb{E}\left[\sum_{i=1}^{n_{b}}Z_{i}^{2}(h)\bar{Z}_{n_{b}}^{2}(h)\right]+n_{b}^{2}\mathbb{E}\left[\bar{Z}_{n_{b},h}^{4}\right]$
$=\mathbb{E}\left[\sum_{i=1}^{n_{b}}Z_{i}^{2}(h)\right]^{2}-2n_{b}^{2}\mathbb{E}\left[Z_{1}^{2}(h)\bar{Z}_{n_{b}}^{2}(h)\right]+n_{b}^{2}\mathbb{E}\left[Z_{1}(h)\bar{Z}_{n_{b}}^{3}(h)\right]$
$=\mathbb{E}\left[\sum_{i=1}^{n_{b}}Z_{i}^{2}(h)\right]^{2}-2\mathbb{E}\left[Z_{1}^{2}(h)\left(\sum_{i=1}^{n_{b}}Z_{i}(h)\right)^{2}\right]+\frac{1}{n_{b}}\mathbb{E}\left[Z_{1}(h)\left(\sum_{i=1}^{n_{b}}Z_{i}(h)\right)^{3}\right],$ (49)
where the fourth equality holds because for any $1\leq i\leq n_{b}$, we have
$\mathbb{E}\left[Z_{i}^{2}(h)\bar{Z}_{n_{b}}^{2}(h)\right]=\mathbb{E}\left[Z_{1}^{2}(h)\bar{Z}_{n_{b}}^{2}(h)\right],\quad\mathbb{E}\left[Z_{i}(h)\bar{Z}_{n_{b}}^{3}(h)\right]=\mathbb{E}\left[Z_{1}(h)\bar{Z}_{n_{b}}^{3}(h)\right].$
In the following, we study the three terms on RHS of (49) separately.
The first term on the RHS of (49) $=\mathbb{E}\left[\sum_{i=1}^{n_{b}}Z_{i}^{4}(h)+2\sum_{i<j}Z_{i}^{2}(h)Z_{j}^{2}(h)\right]$
$=n_{b}\mathbb{E}[Z_{1}(h)]^{4}+n_{b}(n_{b}-1)\mathbb{E}[Z_{1}(h)]^{2}\mathbb{E}[Z_{2}(h)]^{2}$
$=n_{b}\nu_{4}(h)+n_{b}(n_{b}-1)\frac{\sigma^{4}(\theta_{0})+o(1)}{4h^{4}},$ (50)
where $\nu_{4}(h)=\mathbb{E}[Z_{1}(h)]^{4}$ and $\mathbb{E}[Z_{1}(h)]^{2}$ and $\mathbb{E}[Z_{2}(h)]^{2}$ are equal to $\mathrm{Var}[Z_{1}(h)]$ which can be found in (46). Then, we calculate the second term on the RHS of (49) as
The second term on the RHS of (49)
$=$ $-2\mathbb{E}\left[Z_{1}^{2}(h)\left(\sum_{i=1}^{n_{b}}Z_{i}^{2}(h)+2\sum_{i<j}Z_{i}(h)Z_{j}(h)\right)\right]$
$=$ $-2\mathbb{E}[Z_{1}(h)]^{4}-2\mathbb{E}\left[Z_{1}^{2}(h)\sum_{i>1}Z_{i}^{2}(h)\right]-4\mathbb{E}\left[Z_{1}^{2}(h)\sum_{i<j}Z_{i}(h)Z_{j}(h)\right]$
$=$ $-2\nu_{4}(h)-2(n_{b}-1)\mathbb{E}[Z_{1}^{2}(h)]\mathbb{E}[Z_{2}^{2}(h)]-4\mathbb{E}\left[Z_{1}^{3}(h)\sum_{j>1}Z_{j}(h)+Z_{1}^{2}(h)\sum_{1<i<j}Z_{i}(h)Z_{j}(h)\right]$
$=$ $-2\nu_{4}(h)-2(n_{b}-1)\frac{\sigma^{4}(\theta_{0})+o(1)}{4h^{4}},$ (51)
where the last equality is because for any $j>1$, $\mathbb{E}[Z_{1}^{3}(h)Z_{j}(h)]=\mathbb{E}[Z_{1}^{3}(h)]\mathbb{E}[Z_{j}(h)]=0$, and for any $1<i<j$, $\mathbb{E}[Z_{1}^{2}(h)Z_{i}(h)Z_{j}(h)]=\mathbb{E}[Z_{1}^{2}(h)]\mathbb{E}[Z_{i}(h)]\mathbb{E}[Z_{j}(h)]=0$.
The third term on the RHS of (49)
$=$ $\frac{1}{n_{b}}\mathbb{E}\left[Z_{1}(h)\left(\sum_{i=1}^{n_{b}}Z_{i}^{3}(h)+\sum_{i\neq j}3Z_{i}^{2}(h)Z_{j}(h)+\sum_{i<j<k}6Z_{i}(h)Z_{j}(h)Z_{k}(h)\right)\right]$
$=$ $\frac{1}{n_{b}}\mathbb{E}\left[Z_{1}^{4}(h)+Z_{1}(h)\sum_{i>1}Z_{i}^{3}(h)+\sum_{j>1}3Z_{1}^{3}(h)Z_{j}(h)+\sum_{i>1}3Z_{1}^{2}(h)Z_{i}^{2}(h)+\sum_{i,j>1,i\neq j}3Z_{1}(h)Z_{i}^{2}(h)Z_{j}(h)\right.$
$+\left.\sum_{1<j<k}6Z_{1}^{2}(h)Z_{j}(h)Z_{k}(h)+\sum_{1<i<j<k}6Z_{1}(h)Z_{i}(h)Z_{j}(h)Z_{k}(h)\right]$
$=$ $\frac{1}{n_{b}}\mathbb{E}\left[Z_{1}^{4}(h)+\sum_{i>1}3Z_{1}^{2}(h)Z_{i}^{2}(h)\right]=\frac{\nu_{4}(h)}{n_{b}}+\frac{3(n_{b}-1)}{n_{b}}\frac{\sigma^{4}(\theta_{0})+o(1)}{4h^{4}},$ (52)
where the third equality is because for any $i\neq j$, $\mathbb{E}[Z_{i}(h)Z_{j}^{3}(h)]=\mathbb{E}[Z_{i}(h)]\mathbb{E}[Z_{j}^{3}(h)]=0$, and for any $i\neq j,i\neq k,j\neq k$, $\mathbb{E}[Z_{i}^{2}(h)Z_{j}(h)Z_{k}(h)]=\mathbb{E}[Z_{i}^{2}(h)]\mathbb{E}[Z_{j}(h)]\mathbb{E}[Z_{k}(h)]=0,\mathbb{E}[Z_{1}(h)Z_{i}(h)Z_{j}(h)Z_{k}(h)]=\mathbb{E}[Z_{1}(h)]\mathbb{E}[Z_{i}(h)]\mathbb{E}[Z_{j}(h)]\mathbb{E}[Z_{k}(h)]=0$. Combine (49), (50), (51) and (52), leading to
$(n_{b}-1)^{2}\mathbb{E}[S^{2}]^{2}=\frac{(n_{b}-1)^{2}}{n_{b}}\nu_{4}(h)+\frac{(n_{b}^{2}-2n_{b}+3)(n_{b}-1)}{n_{b}}\frac{\sigma^{4}(\theta_{0})+o(1)}{4h^{4}}.$
Therefore,
$\mathbb{E}[S^{2}]^{2}=\frac{\nu_{4}(h)}{n_{b}}+\frac{n_{b}^{2}-2n_{b}+3}{n_{b}(n_{b}-1)}\frac{\sigma^{4}(\theta_{0})+o(1)}{4h^{4}},$
and (48) is
$\mathrm{Var}[\xi(h)]=\frac{\nu_{4}(h)}{n_{b}}-\frac{n_{b}-3}{n_{b}(n_{b}-1)}\frac{\sigma^{4}(\theta_{0})+o(1)}{4h^{4}}.$ (53)
According to (45) and (47), we have
$\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]=\frac{(n_{b}-1)(\sigma^{2}(\theta_{0})+o(1))}{2n_{b}^{2}h^{2}}+\phi(h),$
where $\phi(h)=(n_{b}-1)\xi(h)/n_{b}^{2}$ and
$\mathrm{Var}[\phi(h)]$ $=\frac{(n_{b}-1)^{2}}{n_{b}^{4}}\left(\frac{\nu_{4}(h)}{n_{b}}-\frac{n_{b}-3}{n_{b}(n_{b}-1)}\frac{\sigma^{4}(\theta_{0})+o(1)}{4h^{4}}\right)$
$=\frac{(n_{b}-1)^{2}}{n_{b}^{4}h^{4}}\left(\frac{\nu_{4}+o(1)}{n_{b}}-\frac{n_{b}-3}{n_{b}(n_{b}-1)}\frac{\sigma^{4}(\theta_{0})+o(1)}{4}\right),$
where the second equality is because
$\nu_{4}(h)=\mathbb{E}[Z_{1}(h)]^{4}=\frac{\mathbb{E}[hZ_{1}(h)]^{4}}{h^{4}}=\frac{\nu_{4}+o(1)}{h^{4}}.$
The proof is complete.
A.2 Proof of Theorem 2
Using the least squares method, we get
$\widehat{\bm{\beta}}_{e}$ $=\left(\bm{X}_{e}^{\top}\bm{X}_{e}\right)^{-1}\bm{X}_{e}^{\top}\bm{Y}_{e}.$
Note that
$\bm{\beta}_{e}$ $=\left(\bm{X}_{e}^{\top}\bm{X}_{e}\right)^{-1}\bm{X}_{e}^{\top}\bm{X}_{e}\bm{\beta}_{e}.$
Then, according to (15), we have
$\widehat{\bm{\beta}}_{e}-\bm{\beta}_{e}$ $=\left(\bm{X}_{e}^{\top}\bm{X}_{e}\right)^{-1}\bm{X}_{e}^{\top}(\bm{Y}_{e}-\bm{X}_{e}\bm{\beta}_{e})=\left(\bm{X}_{e}^{\top}\bm{X}_{e}\right)^{-1}\bm{X}_{e}^{\top}\mathcal{E}_{e}$
$=\left(\left[\begin{array}[]{lll}1&...&1\\ h_{1}^{2}&...&h_{K}^{2}\end{array}\right]\left[\begin{array}[]{lll}1&...&1\\ h_{1}^{2}&...&h_{K}^{2}\end{array}\right]^{\top}\right)^{-1}\left[\begin{array}[]{lll}1&...&1\\ h_{1}^{2}&...&h_{K}^{2}\end{array}\right]\left[\begin{array}[]{c}Dh_{1}^{4}+o(h_{1}^{4})+\bar{Z}_{n_{b},h_{1}}\\ \vdots\\ Dh_{K}^{4}+o(h_{K}^{4})+\bar{Z}_{n_{b},h_{K}}\end{array}\right]$
$=\left[\begin{array}[]{cc}K&\sum_{k=1}^{K}h_{k}^{2}\\ \sum_{k=1}^{K}h_{k}^{2}&\sum_{k=1}^{K}h_{k}^{4}\end{array}\right]^{-1}\left[\begin{array}[]{c}\sum_{k=1}^{K}\left(Dh_{k}^{4}+o(h_{k}^{4})+\bar{Z}_{n_{b},h_{k}}\right)\\ \sum_{k=1}^{K}\left(Dh_{k}^{6}+o(h_{k}^{4})+h_{k}^{2}\bar{Z}_{n_{b},h_{k}}\right)\end{array}\right]$
$=\frac{\left[\begin{array}[]{cc}\sum_{k=1}^{K}h_{k}^{4}&-\sum_{k=1}^{K}h_{k}^{2}\\ -\sum_{k=1}^{K}h_{k}^{2}&K\end{array}\right]\left[\begin{array}[]{c}\sum_{k=1}^{K}\left(Dh_{k}^{4}+o(h_{k}^{4})+\bar{Z}_{n_{b},h_{k}}\right)\\ \sum_{k=1}^{K}\left(Dh_{k}^{6}+o(h_{k}^{6})+h_{k}^{2}\bar{Z}_{n_{b},h_{k}}\right)\end{array}\right]}{K\sum_{k=1}^{K}h_{k}^{4}-\left(\sum_{k=1}^{K}h_{k}^{2}\right)^{2}}.$
Therefore,
$\widehat{B}-B=\frac{K\sum_{k=1}^{K}\left(Dh_{k}^{6}+o(h_{k}^{6})+h_{k}^{2}\bar{Z}_{n_{b},h_{k}}\right)-\sum_{k=1}^{K}h_{k}^{2}\sum_{k=1}^{K}\left(Dh_{k}^{4}+o(h_{k}^{4})+\bar{Z}_{n_{b},h_{k}}\right)}{K\sum_{k=1}^{K}h_{k}^{4}-\left(\sum_{k=1}^{K}h_{k}^{2}\right)^{2}}.$
We divide $\widehat{B}-B$ into two parts, i.e.,
$\widehat{B}-B=$ $\frac{K\sum_{k=1}^{K}\left(Dh_{k}^{6}+o(h_{k}^{6})\right)-\sum_{k=1}^{K}h_{k}^{2}\sum_{k=1}^{K}\left(Dh_{k}^{4}+o(h_{k}^{4})\right)}{K\sum_{k=1}^{K}h_{k}^{4}-\left(\sum_{k=1}^{K}h_{k}^{2}\right)^{2}}$ (54)
$+\frac{K\sum_{k=1}^{K}h_{k}^{2}\bar{Z}_{n_{b},h_{k}}-\sum_{k=1}^{K}h_{k}^{2}\sum_{k=1}^{K}\bar{Z}_{n_{b},h_{k}}}{K\sum_{k=1}^{K}h_{k}^{4}-\left(\sum_{k=1}^{K}h_{k}^{2}\right)^{2}}.$ (55)
Note that (54) is the non-random term, while (55) is the random term whose variance is the same as that of $\widehat{B}$. We now separately calculate the bias, variance, and then the MSE of $\widehat{B}$.
For any $h_{k}$, $\mathbb{E}\left[\bar{Z}_{n_{b},h_{k}}\right]=0$, so the expectation of (55) is 0. Furthermore, the bias of $\widehat{B}$ is
$\mbox{Bias}\left[\widehat{B}\right]=\mathbb{E}\left[\widehat{B}-B\right]$ $=\frac{K\sum_{k=1}^{K}\left(Dh_{k}^{6}+o(h_{k}^{6})\right)-\sum_{k=1}^{K}h_{k}^{2}\sum_{k=1}^{K}\left(Dh_{k}^{4}+o(h_{k}^{4})\right)}{K\sum_{k=1}^{K}h_{k}^{4}-\left(\sum_{k=1}^{K}h_{k}^{2}\right)^{2}}$
$=\frac{K\sum_{k=1}^{K}Dc_{k}^{6}-\sum_{k=1}^{K}c_{k}^{2}\sum_{k=1}^{K}Dc_{k}^{4}}{K\sum_{k=1}^{K}c_{k}^{4}-\left(\sum_{k=1}^{K}c_{k}^{2}\right)^{2}}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right)$
$=H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right),$ (56)
where
$H_{K}=\frac{K\sum_{k=1}^{K}Dc_{k}^{6}-\sum_{k=1}^{K}c_{k}^{2}\sum_{k=1}^{K}Dc_{k}^{4}}{K\sum_{k=1}^{K}c_{k}^{4}-\left(\sum_{k=1}^{K}c_{k}^{2}\right)^{2}}.$
For the variance of $\widehat{B}$, by (55),
$\mathrm{Var}\left[\widehat{B}\right]\left(K\sum_{k=1}^{K}h_{k}^{4}-\left(\sum_{k=1}^{K}h_{k}^{2}\right)^{2}\right)^{2}=\mathrm{Var}\left[K\sum_{k=1}^{K}h_{k}^{2}\bar{Z}_{n_{b},h_{k}}-\sum_{k=1}^{K}h_{k}^{2}\sum_{k=1}^{K}\bar{Z}_{n_{b},h_{k}}\right]$
$=$ $K^{2}\sum_{k=1}^{K}h_{k}^{4}\mathrm{Var}\left[\bar{Z}_{n_{b},h_{k}}\right]-2K\sum_{k=1}^{K}h_{k}^{2}\sum_{k=1}^{K}h_{k}^{2}\mathrm{Var}\left[\bar{Z}_{n_{b},h_{k}}\right]+\left(\sum_{k=1}^{K}h_{k}^{2}\right)^{2}\sum_{k=1}^{K}\mathrm{Var}\left[\bar{Z}_{n_{b},h_{k}}\right]$
$=$ $\frac{\sigma^{2}(\theta_{0})+o(1)}{2n_{b}}\left(K^{2}\sum_{k=1}^{K}h_{k}^{2}-2K^{2}\sum_{k=1}^{K}h_{k}^{2}+\left(\sum_{k=1}^{K}h_{k}^{2}\right)^{2}\sum_{k=1}^{K}\frac{1}{h_{k}^{2}}\right)$
$=$ $\frac{\sigma^{2}(\theta_{0})+o(1)}{2n_{b}^{1+6\gamma}}\left(-K^{2}\sum_{k=1}^{K}c_{k}^{2}+\left(\sum_{k=1}^{K}c_{k}^{2}\right)^{2}\sum_{k=1}^{K}\frac{1}{c_{k}^{2}}\right).$
Furthermore, the variance of $\widehat{B}$ is
$\mathrm{Var}\left[\widehat{B}\right]=V_{K}\frac{\sigma^{2}(\theta_{0})+o(1)}{2n_{b}^{1+6\gamma}},$ (57)
where
$V_{K}=\frac{-K^{2}\sum_{k=1}^{K}c_{k}^{2}+\left(\sum_{k=1}^{K}c_{k}^{2}\right)^{2}\sum_{k=1}^{K}1/c_{k}^{2}}{\left(K\sum_{k=1}^{K}c_{k}^{4}-\left(\sum_{k=1}^{K}c_{k}^{2}\right)^{2}\right)^{2}}.$
Combining (A.2) and (57) gives the MSE of $\widehat{B}$.
$\mbox{MSE}\left[\widehat{B}\right]=\mbox{Bias}^{2}\left[\widehat{B}\right]+\mathrm{Var}\left[\widehat{B}\right]=H_{K}^{2}n_{b}^{4\gamma}+o\left(n_{b}^{4\gamma}\right)+V_{K}\frac{\sigma^{2}(\theta_{0})+o(1)}{2n_{b}^{1+6\gamma}}.$ (58)
From (58), when $-1/6<\gamma<0$, we have $\mbox{MSE}\left[\widehat{B}\right]\to 0$ as $n_{b}\to\infty$. Therefore, according to Chebyshev’s inequality, for any $\epsilon>0$, as $n_{b}\to\infty$
$\mathbb{P}\left(\left|\widehat{B}-B\right|\geq\epsilon\right)\leq\frac{\mathbb{E}\left[\widehat{B}-B\right]^{2}}{\epsilon^{2}}=\frac{\mbox{MSE}\left[\widehat{B}\right]}{\epsilon^{2}}\to 0.$
So $\widehat{B}\stackrel{{\scriptstyle p}}{{\longrightarrow}}B$. The proof is complete.
A.3 Asymptotic property of $\widehat{\alpha}^{\prime}(\theta_{0})$
Corollary 6.
Under the same conditions as in Theorem 2, we have
$\mathbb{E}\left[\widehat{\alpha}^{\prime}(\theta_{0})-\alpha^{\prime}(\theta_{0})\right]=\widetilde{H}_{K}n_{b}^{4\gamma}+o\left(n_{b}^{4\gamma}\right),\quad\mathrm{Var}\left[\widehat{\alpha}^{\prime}(\theta_{0})\right]=\widetilde{V}_{K}\frac{\sigma^{2}(\theta_{0})+o(1)}{2n_{b}^{1+2\gamma}},$ (59)
where $\widetilde{H}_{K}$ and $\widetilde{V}_{K}$ are constants depending on $c_{1},...,c_{K}$. Specifically,
$\widetilde{H}_{K}$ $=\frac{\sum_{k=1}^{K}c_{k}^{4}\sum_{k=1}^{K}Dc_{k}^{4}-\sum_{k=1}^{K}c_{k}^{2}\sum_{k=1}^{K}Dc_{k}^{6}}{K\sum_{k=1}^{K}c_{k}^{4}-\left(\sum_{k=1}^{K}c_{k}^{2}\right)^{2}},$
$\widetilde{V}_{K}$ $=\frac{\left(\sum_{k=1}^{K}c_{k}^{2}\right)^{3}-2K\sum_{k=1}^{K}c_{k}^{4}\sum_{k=1}^{K}c_{k}^{2}+\left(\sum_{k=1}^{K}c_{k}^{4}\right)^{2}\sum_{k=1}^{K}1/c_{k}^{2}}{\left({K\sum_{k=1}^{K}c_{k}^{4}-\left(\sum_{k=1}^{K}c_{k}^{2}\right)^{2}}\right)^{2}}.$
In addition, if $-1/6<\gamma<0$, then $\widehat{\alpha}^{\prime}(\theta_{0})\stackrel{{\scriptstyle p}}{{\longrightarrow}}\alpha^{\prime}(\theta_{0})$.
The proof of Corollary 6 is identical to that of Theorem 2. (59) provides the convergence rates for the bias and the variance of $\widehat{\alpha}^{\prime}(\theta_{0})$, which are $O\left(n_{b}^{4\gamma}\right)$ and $O\left(n_{b}^{-1-2\gamma}\right)$, respectively. Additionally, when $c_{1},...,c_{K}$ are generated from a proper distribution $\mathcal{P}_{0}$, the convergence rates of the bias and the variance of $\widehat{\alpha}^{\prime}(\theta_{0})$ are $O\left(n_{b}^{4\gamma}\right)$ and $O\left(n_{b}^{8\gamma}\right)+O\left(n_{b}^{-1-2\gamma}\right)$, respectively. Theorem 2 and Corollary 6 demonstrate that when estimating $B$, we also obtain a roughly estimate of $\alpha^{\prime}(\theta_{0})$, which is denoted by $\widehat{\alpha}^{\prime}(\theta_{0})$. However, directly using $\widehat{\alpha}^{\prime}(\theta_{0})$ is not advisable since it bases on a small sample size $Kn_{b}$. Nevertheless, it is evident that $\widehat{\alpha}^{\prime}(\theta_{0})$ contains valuable gradient information. To prevent the loss of the information, in Section 4, we will propose a sample-recycling-based estimator which contains $\widehat{\alpha}^{\prime}(\theta_{0})$.
A.4 Analysis of (21)
Under Assumption 4, the left hand side (LHS) of (46) is
$\mathbb{E}[S^{2}]$ $=\mathrm{Var}[Z_{1}(h)]=\mathrm{Var}[\Delta_{1}(h)]=\frac{\sigma^{2}(\theta_{0}+h)+\sigma^{2}(\theta_{0}-h)}{4h^{2}}$
$=\frac{\left(\sigma(\theta_{0})+\sigma^{\prime}(\theta_{0})h+o(h)\right)^{2}+\left(\sigma(\theta_{0})-\sigma^{\prime}(\theta_{0})h+o(h)\right)^{2}}{4h^{2}}$
$=\frac{\sigma^{2}(\theta_{0})+\sigma^{\prime}(\theta_{0})^{2}h^{2}+o(h^{2})}{2h^{2}},$
where the fourth equality is due to the Taylor expansion. Then, following the proof of Theorem 1, (21) is obtained.
A.5 Proof of Theorem 3
Using the least squares method, we get
$\widehat{\bm{\beta}}_{v}$ $=\left(\bm{X}_{v}^{\top}\bm{X}_{v}\right)^{-1}\bm{X}_{v}^{\top}\bm{Y}_{v}.$
Note that
$\bm{\beta}_{v}$ $=\left(\bm{X}_{v}^{\top}\bm{X}_{v}\right)^{-1}\bm{X}_{v}^{\top}\bm{X}_{v}\bm{\beta}_{v}.$
Then, according to (22), we have
$\widehat{\bm{\beta}}_{v}-\bm{\beta}_{v}$ $=\left(\bm{X}_{v}^{\top}\bm{X}_{v}\right)^{-1}\bm{X}_{v}^{\top}(\bm{Y}_{v}-\bm{X}_{v}\bm{\beta}_{v})=\left(\bm{X}_{v}^{\top}\bm{X}_{v}\right)^{-1}\bm{X}_{v}^{\top}\mathcal{E}_{v}$
$=\left(\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}}{4n_{b}^{4}h_{k}^{4}}\right)^{-1}\sum_{k=1}^{K}\left(\frac{(n_{b}-1)^{2}\left[\sigma^{\prime}(\theta_{0})^{2}h_{k}^{2}+o(h_{k}^{2})\right]}{4n_{b}^{4}h_{k}^{4}}+\frac{n_{b}-1}{2n_{b}^{2}h_{k}^{2}}\phi(h_{k})\right).$
We divide $\widehat{\bm{\beta}}_{v}-\bm{\beta}_{v}$ into two parts, i.e.,
$\widehat{\bm{\beta}}_{v}-\bm{\beta}_{v}=$ $\left(\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}}{4n_{b}^{4}h_{k}^{4}}\right)^{-1}\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}\left[\sigma^{\prime}(\theta_{0})^{2}h_{k}^{2}+o(h_{k}^{2})\right]}{4n_{b}^{4}h_{k}^{4}}$ (60)
$+\left(\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}}{4n_{b}^{4}h_{k}^{4}}\right)^{-1}\sum_{k=1}^{K}\frac{n_{b}-1}{2n_{b}^{2}h_{k}^{2}}\phi(h_{k}).$ (61)
Note that (60) is the non-random term, while (61) is the random term whose variance is the same as that of $\widehat{\bm{\beta}}_{v}$. We now separately calculate the bias, variance, and then the MSE of $\widehat{\bm{\beta}}_{v}$.
For any $h_{k}$, $\mathbb{E}\left[\phi(h_{k})\right]=0$, so the expectation of (61) is 0. Furthermore, the bias of $\widehat{\bm{\beta}}_{v}$ is
$\mbox{Bias}\left[\widehat{\bm{\beta}}_{v}\right]=\mathbb{E}\left[\widehat{\bm{\beta}}_{v}-\bm{\beta}_{v}\right]$ $=\left(\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}}{4n_{b}^{4}h_{k}^{4}}\right)^{-1}\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}\left[\sigma^{\prime}(\theta_{0})^{2}h_{k}^{2}+o(h_{k}^{2})\right]}{4n_{b}^{4}h_{k}^{4}}$
$=\left(\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}}{4n_{b}^{4}c_{k}^{4}n_{b}^{4\gamma}}\right)^{-1}\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}\left[\sigma^{\prime}(\theta_{0})^{2}c_{k}^{2}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right)\right]}{4n_{b}^{4}c_{k}^{4}n_{b}^{4\gamma}}$
$=\frac{4n_{b}^{4+4\gamma}}{(n_{b}-1)^{2}}\left(\sum_{k=1}^{K}\frac{1}{c_{k}^{4}}\right)^{-1}\frac{\sigma^{\prime}(\theta_{0})^{2}(n_{b}-1)^{2}n_{b}^{2\gamma}}{4n_{b}^{4+4\gamma}}\sum_{k=1}^{K}\frac{1+o(1)}{c_{k}^{2}}$
$=\widehat{H}_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right),$ (62)
where
$\widehat{H}_{K}=\sigma^{\prime}(\theta_{0})^{2}\left(\sum_{k=1}^{K}\frac{1}{c_{k}^{4}}\right)^{-1}\sum_{k=1}^{K}\frac{1}{c_{k}^{2}}.$
The variance of $\widehat{\bm{\beta}}_{v}$ is
$\mathrm{Var}\left[\widehat{\bm{\beta}}_{v}\right]$ $=\left(\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}}{4n_{b}^{4}h_{k}^{4}}\right)^{-2}\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}}{4n_{b}^{4}h_{k}^{4}}\mathrm{Var}[\phi(h_{k})]$
$=\left(\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}}{4n_{b}^{4}c_{k}^{4}n_{b}^{4\gamma}}\right)^{-2}\sum_{k=1}^{K}\frac{(n_{b}-1)^{2}}{4n_{b}^{4}c_{k}^{4}n_{b}^{4\gamma}}\mathrm{Var}[\phi(h_{k})]$
$=\frac{16n_{b}^{8+8\gamma}}{(n_{b}-1)^{4}}\left(\sum_{k=1}^{K}\frac{1}{c_{k}^{4}}\right)^{-2}\frac{(n_{b}-1)^{2}}{4n_{b}^{4+4\gamma}}\sum_{k=1}^{K}\frac{\mathrm{Var}[\phi(h_{k})]}{c_{k}^{4}}$
$=\frac{4n_{b}^{4+4\gamma}}{(n_{b}-1)^{2}}\left(\sum_{k=1}^{K}\frac{1}{c_{k}^{4}}\right)^{-2}\sum_{k=1}^{K}\frac{\mathrm{Var}[\phi(h_{k})]}{c_{k}^{4}}$
$=\frac{4n_{b}^{4+4\gamma}}{(n_{b}-1)^{2}}\left(\sum_{k=1}^{K}\frac{1}{c_{k}^{4}}\right)^{-2}\sum_{k=1}^{K}\frac{1}{c_{k}^{4}}\frac{(n_{b}-1)^{2}}{n_{b}^{4}h_{k}^{4}}\left(\frac{\nu_{4}+o(1)}{n_{b}}-\frac{n_{b}-3}{n_{b}(n_{b}-1)}\frac{\sigma^{4}(\theta_{0})+o(1)}{4}\right)$
$=\frac{4n_{b}^{4+4\gamma}}{(n_{b}-1)^{2}}\left(\sum_{k=1}^{K}\frac{1}{c_{k}^{4}}\right)^{-2}\sum_{k=1}^{K}\frac{1}{c_{k}^{4}}\frac{(n_{b}-1)^{2}}{n_{b}^{4}c_{k}^{4}n_{b}^{4\gamma}}\left(\frac{\nu_{4}+o(1)}{n_{b}}-\frac{n_{b}-3}{n_{b}(n_{b}-1)}\frac{\sigma^{4}(\theta_{0})+o(1)}{4}\right)$
$=\widehat{V}_{K}\frac{4\nu_{4}(n_{b}-1)-\sigma^{4}(\theta_{0})(n_{b}-3)}{n_{b}(n_{b}-1)}+o\left(\frac{1}{n_{b}}\right),$ (63)
where the fifth equality is from Theorem 1 and
$\widehat{V}_{K}=\left(\sum_{k=1}^{K}\frac{1}{c_{k}^{4}}\right)^{-2}\sum_{k=1}^{K}\frac{1}{c_{k}^{8}}.$
Combining (A.5) and (A.5) gives the MSE of $\widehat{\bm{\beta}}_{v}$.
$\mbox{MSE}\left[\widehat{\bm{\beta}}_{v}\right]$ $=\mbox{Bias}^{2}\left[\widehat{\bm{\beta}}_{v}\right]+\mathrm{Var}\left[\widehat{\bm{\beta}}_{v}\right]$
$=\widehat{H}_{K}^{2}n_{b}^{4\gamma}+o\left(n_{b}^{4\gamma}\right)+\widehat{V}_{K}\frac{4\nu_{4}(n_{b}-1)-\sigma^{4}(\theta_{0})(n_{b}-3)}{n_{b}(n_{b}-1)}+o\left(\frac{1}{n_{b}}\right)$ (64)
From (A.5), when $\gamma<0$, we have $\mbox{MSE}\left[\widehat{\bm{\beta}}_{v}\right]\to 0$ as $n_{b}\to\infty$. Therefore, according to Chebyshev’s inequality, for any $\epsilon>0$, as $n_{b}\to\infty$
$\mathbb{P}\left(\left|\widehat{\bm{\beta}}_{v}-\bm{\beta}_{v}\right|\geq\epsilon\right)\leq\frac{\mathbb{E}\left[\widehat{\bm{\beta}}_{v}-\bm{\beta}_{v}\right]^{2}}{\epsilon^{2}}=\frac{\mbox{MSE}\left[\widehat{\bm{\beta}}_{v}\right]}{\epsilon^{2}}\to 0.$
The proof is complete.
Appendix B B Discussions And Proofs in Section 4
B.1 Further explanations on (27)
Note that the continuous mapping theorem implies that $\left(\frac{\widehat{\bm{\beta}}_{v}}{\widehat{B}^{2}}\right)^{1/3}$ and $\left(\frac{\widehat{B}^{2}}{\widehat{\bm{\beta}}_{v}}\right)^{1/3}$ converge to $\left(\frac{\bm{\beta}_{v}}{B^{2}}\right)^{1/3}$ and $\left(\frac{B^{2}}{\bm{\beta}_{v}}\right)^{1/3}$ in probability, respectively. Note that the expectation and limit operators may not be exchangeable because 0 exists at the support set of the regression estimators. However, we don’t have to be confused about this issue. Take $\widehat{B}$ for example, assume that $B$ is bounded below away from 0. Specifically, we assume that there exists a positive number $\epsilon$ such that $B>\epsilon>0$. We can solve (14) under the constraint $\widehat{B}>\epsilon$. In others words, we denote a new estimator $\widetilde{B}=\epsilon+\max{\left\{\widehat{B}-\epsilon,0\right\}}$ and replace $\widehat{B}$ with $\widetilde{B}$. Lemma 7 and Proposition 8 indicate that the asymptotic property of $\widetilde{B}$ is identical to that of $\widehat{B}$. As a result, the dominated convergence theorem can be used to obtain the leading terms of (27).
Lemma 7.
Suppose that Assumption 3 holds and assume that $f(\cdot)$ is third continuously differentiable and has a uniformly bounded third derivative. For any $k=1,...,K(K\geq 2)$, denote $h_{k}=c_{k}n_{b}^{\gamma}(c_{k}\neq 0,\gamma<0)$ and for any $j\neq k$, let $c_{j}\neq c_{k}$. Then,
$\mathbb{E}\left[f\left(\widehat{B}\right)\right]$ $=f(B)+f^{\prime}(B)H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right),$ (65)
$\mathrm{Var}\left[f\left(\widehat{B}\right)\right]=f(B)f^{\prime\prime}(B)H_{K}^{2}n_{b}^{4\gamma}+$ $\left(f^{\prime}(B)^{2}+f(B)f^{\prime\prime}(B)\right)V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right),$ (66)
where $H_{K}$ and $V_{K}$ are defined as in Theorem 2.
Proof: We prove (65) and (66), respectively.
Proof of (65): By Taylor expansion, we have
$f\left(\widehat{B}\right)=f(B)+f^{\prime}(B)\left(\widehat{B}-B\right)+\frac{f^{\prime\prime}\left(\breve{B}\right)}{2}\left(\widehat{B}-B\right)^{2},$
where $\breve{B}$ is a random variable between $B$ and $\widehat{B}$. Therefore,
$\mathbb{E}\left[f\left(\widehat{B}\right)\right]=f(B)+f^{\prime}(B)\mathbb{E}\left[\widehat{B}-B\right]+o\left(n_{b}^{2\gamma}\right)=f(B)+f^{\prime}(B)H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right),$ (67)
where $H_{K}$ is defined as in Theorem 2.
Proof of (66): Note that
$\mathrm{Var}\left[f\left(\widehat{B}\right)\right]=\mathbb{E}\left[f^{2}\left(\widehat{B}\right)\right]-\mathbb{E}^{2}\left[f\left(\widehat{B}\right)\right].$ (68)
Our task now is to calculate $\mathbb{E}\left[f^{2}\left(\widehat{B}\right)\right]$. According to Taylor expansion,
$f^{2}\left(\widehat{B}\right)=$ $f^{2}(B)+2f(B)f^{\prime}(B)\left(\widehat{B}-B\right)+\left(f^{\prime}(B)^{2}+f(B)f^{\prime\prime}(B)\right)\left(\widehat{B}-B\right)^{2}$
$+\frac{1}{3}\left(3f^{\prime}\left(\mathring{B}\right)f^{\prime\prime}\left(\mathring{B}\right)+f\left(\mathring{B}\right)f^{\prime\prime\prime}\left(\mathring{B}\right)\right)\left(\widehat{B}-B\right)^{3},$
where $\mathring{B}$ is a random variable between $B$ and $\widehat{B}$. Therefore,
$\mathbb{E}\left[f^{2}\left(\widehat{B}\right)\right]=$ $f^{2}(B)+2f(B)f^{\prime}(B)\mathbb{E}\left[\widehat{B}-B\right]+\left(f^{\prime}(B)^{2}+f(B)f^{\prime\prime}(B)\right)\mathbb{E}\left[\widehat{B}-B\right]^{2}$
$+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right),$
$=$ $f^{2}(B)+2f(B)f^{\prime}(B)H_{K}n_{b}^{2\gamma}+\left(f^{\prime}(B)^{2}+f(B)f^{\prime\prime}(B)\right)\left(H_{K}^{2}n_{b}^{4\gamma}+V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}\right)$
$+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right),$ (69)
where $V_{K}$ is defined as in Theorem 2.
Combining (67), (68) and (B.1) gives that
$\mathrm{Var}\left[f\left(\widehat{B}\right)\right]=f(B)f^{\prime\prime}(B)H_{K}^{2}n_{b}^{4\gamma}+$ $\left(f^{\prime}(B)^{2}+f(B)f^{\prime\prime}(B)\right)V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right).$
The proof is complete.
Proposition 8.
Suppose that Assumption 3 holds and assume that $B$ is bounded below away from 0. For any $k=1,...,K(K\geq 2)$, denote $h_{k}=c_{k}n_{b}^{\gamma}(c_{k}\neq 0,\gamma<0)$ and for any $j\neq k$, let $c_{j}\neq c_{k}$. For $0<\epsilon<B$, define $\widetilde{B}=\epsilon+\max{\left\{\widehat{B}-\epsilon,0\right\}}$. Then,
$\mathbb{E}\left[\widetilde{B}\right]-B=H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right),\quad\mathrm{Var}\left[\widetilde{B}\right]=V_{K}\frac{\sigma^{2}(\theta_{0})+o(1)}{2n_{b}^{1+6\gamma}},$ (70)
where $H_{K}$ and $V_{K}$ are defined as in Theorem 2.
Proof: For a given $\delta>0$, we construct a smooth approximation function of $f(x)=\epsilon+\max{\{x-\epsilon,0\}}$, denoted by $f_{\delta}(x)$, as follows:
$f_{\delta}(x)=\epsilon+\left(\frac{1}{2}(x+\delta-\epsilon)-\frac{\delta}{\pi}\cos\left(\frac{x-\epsilon}{2\delta}\pi\right)\right)\mathbbm{1}_{\{\epsilon-\delta\leq x\leq\epsilon+\delta\}}+(x-\epsilon)\mathbbm{1}_{\{x>\epsilon+\delta\}}.$
It is easily verified that for a fixed $\delta>0$, $f_{\delta}(x)$ is third continuously differentiable and has a uniformly bounded third derivative. Specifically,
$f^{\prime}_{\delta}(x)=$ $\left(\frac{1}{2}+\frac{1}{2}\sin\left(\frac{x-\epsilon}{2\delta}\pi\right)\right)\mathbbm{1}_{\{\epsilon-\delta\leq x\leq\epsilon+\delta\}}+\mathbbm{1}_{\{x>\epsilon+\delta\}},$
$f^{\prime\prime}_{\delta}(x)=$ $\frac{\pi}{4\delta}\cos\left(\frac{x-\epsilon}{2\delta}\pi\right)\mathbbm{1}_{\{\epsilon-\delta\leq x\leq\epsilon+\delta\}},$
$f^{\prime\prime\prime}_{\delta}(x)=$ $-\frac{\pi^{2}}{8\delta^{2}}\sin\left(\frac{x-\epsilon}{2\delta}\pi\right)\mathbbm{1}_{\{\epsilon-\delta\leq x\leq\epsilon+\delta\}}.$
Therefore, according Lemma 7, for a fixed $\delta>0$, we have
$\mathbb{E}\left[f_{\delta}\left(\widehat{B}\right)\right]$ $=f_{\delta}(B)+f_{\delta}^{\prime}(B)H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right),$
$\mathrm{Var}\left[f_{\delta}\left(\widehat{B}\right)\right]=f_{\delta}(B)f_{\delta}^{\prime\prime}(B)H_{K}^{2}n_{b}^{4\gamma}+$ $\left(f_{\delta}^{\prime}(B)^{2}+f_{\delta}(B)f_{\delta}^{\prime\prime}(B)\right)V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right),$
where $H_{K}$ and $V_{K}$ are defined as in Theorem 2. As $\delta\to 0$, by the dominated convergence theorem we get
$\mathbb{E}\left[\widetilde{B}\right]=$ $\mathbb{E}\left[\epsilon+\max{\left\{\widehat{B}-\epsilon,0\right\}}\right]=\mathbb{E}\left[\underset{\delta\to 0}{\lim}f_{\delta}\left(\widehat{B}\right)\right]=\underset{\delta\to 0}{\lim}\mathbb{E}\left[f_{\delta}\left(\widehat{B}\right)\right]$
$=$ $\underset{\delta\to 0}{\lim}f_{\delta}(B)+\underset{\delta\to 0}{\lim}f_{\delta}^{\prime}(B)H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right)$
$=$ $\epsilon+(B-\epsilon)\mathbbm{1}_{\{B>\epsilon\}}+\mathbbm{1}_{\{B>\epsilon\}}H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right)$
$=$ $B+H_{K}n_{b}^{2\gamma}+o\left(n_{b}^{2\gamma}\right),$
and
$\mathrm{Var}\left[\widetilde{B}\right]=$ $\mathrm{Var}\left[\epsilon+\max{\left\{\widehat{B}-\epsilon,0\right\}}\right]=\mathrm{Var}\left[\underset{\delta\to 0}{\lim}f_{\delta}\left(\widehat{B}\right)\right]=\underset{\delta\to 0}{\lim}\mathrm{Var}\left[f_{\delta}\left(\widehat{B}\right)\right]$
$=$ $\underset{\delta\to 0}{\lim}f_{\delta}^{\prime}(B)^{2}V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right)$
$=$ $\mathbbm{1}_{\{B>\epsilon\}}V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right)$
$=$ $V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}+o\left(n_{b}^{4\gamma}\right)+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right),$ (71)
where the fourth equality is since that $B>\epsilon$, there exists $\delta^{\prime}>0$ such that $B>\epsilon+\delta^{\prime}$. Therefore, for any $\delta<\delta^{\prime}$, $f_{\delta}^{\prime\prime}(B)=\dfrac{\pi}{4\delta}\cos\left(\dfrac{B-\epsilon}{2\delta}\pi\right)\mathbbm{1}_{\{\epsilon-\delta\leq B\leq\epsilon+\delta\}}=0$, i.e., $\underset{\delta\to 0}{\lim}f_{\delta}^{\prime\prime}(B)=0$. Similarly, $\underset{\delta\to 0}{\lim}f_{\delta}^{\prime\prime\prime}(B)=0$. As a result, it is obvious that $o\left(n_{b}^{4\gamma}\right)$ in (B.1) is 0. Therefore,
$\mathrm{Var}\left[\widetilde{B}\right]=V_{K}\frac{\sigma^{2}(\theta_{0})}{2n_{b}^{1+6\gamma}}+o\left(\frac{1}{n_{b}^{1+6\gamma}}\right).$
The proof is complete.
B.2 Proof of Theorem 4
From (3), (31) and (32), it follows that
$\widehat{\Delta}_{n,\widehat{h}_{n}}=\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+n_{b}\left(\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\widehat{\Delta}_{n_{b},h_{k}}-\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}\right)+K\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}\right)\right),$
where $\widehat{\bm{\beta}}_{e}=\left(\bm{X}_{e}^{\top}\bm{X}_{e}\right)^{-1}\bm{X}_{e}^{\top}\bm{Y}_{e}$. From (43),
$\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h}=\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}\Delta_{i}(h)=\widehat{\Delta}_{n_{b},h},$
so $\widehat{\bm{\beta}}_{e}=\left(\bm{X}_{e}^{\top}\bm{X}_{e}\right)^{-1}\bm{X}_{e}^{\top}\left[\widehat{\Delta}_{n_{b},h_{1}},...,\widehat{\Delta}_{n_{b},h_{K}}\right]^{\top}$ and we can represent $\widehat{\Delta}_{n,\widehat{h}_{n}}$ in matrix form
$\widehat{\Delta}_{n,\widehat{h}_{n}}=\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+n_{b}\bm{h}^{\top}\bm{P}\bm{\Delta_{h}}+Kn_{b}\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}\right),$
where ${\bm{h}}=[|h_{1}|,...,|h_{K}|]^{\top}$, ${\bm{\Delta}_{h}}=\left[\widehat{\Delta}_{n_{b},h_{1}}\big{/}\left|\widehat{h}_{n}\right|,...,\widehat{\Delta}_{n_{b},h_{K}}\big{/}\left|\widehat{h}_{n}\right|\right]^{\top}$ and ${\bm{P}}=\bm{I}-\bm{X}_{e}(\bm{X}_{e}^{\top}\bm{X}_{e})^{-1}\bm{X}_{e}^{\top}$. Note that (35) is a straightforward corollary of (33) and (34), so we only have to prove (33) and (34) separately.
Proof of (33): According to the linearity of the expectation,
$n\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}=\mathbb{E}\left[\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)\right]+Kn_{b}\mathbb{E}\left[\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}\right]+n_{b}\mathbb{E}\left[\bm{h}^{\top}\bm{P}\bm{\Delta_{h}}\right].$ (72)
We consider the three terms on the RHS of (72) separately. Firstly, we study the first term on the RHS of (72).
[…]
$=$ $\mathbb{E}\left[n_{2}\left(\alpha^{\prime}(\theta_{0})+B\widehat{h}_{n}^{2}+o\left(\widehat{h}_{n}^{2}\right)\right)\right]$
$=$ $n_{2}\left(\mathbb{E}\left[\alpha^{\prime}(\theta_{0})+B\widehat{h}_{n}^{2}\right]+o\left(n^{-1/3}\right)\right),$ (73)
where the second equality is due to (5). As for the third equality, we omit the lengthy discussion on the technical assumptions needed to ensure $\mathbb{E}\left[o\left(\widehat{h}_{n}^{2}\right)\right]=o\left(n^{-1/3}\right)$5. In the following, we keep assuming that the operators $\mathbb{E}[\cdot]$ (or $\mathrm{Var}[\cdot]$) and $o(\cdot)$ (or $o_{p}(\cdot)$) can be exchanged.
Secondly, we consider the second term on the RHS of (72).
The second term on the RHS of (72)
$=$ $Kn_{b}\mathbb{E}\left[\left(1,{\widehat{h}_{n}}^{2}\right){\bm{\beta}}_{e}+\left(1,{\widehat{h}_{n}}^{2}\right)\left(\widehat{\bm{\beta}}_{e}-{\bm{\beta}}_{e}\right)\right]$
$=$ $Kn_{b}\mathbb{E}\left[\alpha^{\prime}(\theta_{0})+B\widehat{h}_{n}^{2}\right]+Kn_{b}\mathbb{E}\left[\widehat{\alpha}^{\prime}(\theta_{0})-\alpha^{\prime}(\theta_{0})\right]+Kn_{b}\mathbb{E}\left[\left(\widehat{B}-B\right)\widehat{h}_{n}^{2}\right]$
$=$ $Kn_{b}\mathbb{E}\left[\alpha^{\prime}(\theta_{0})+B\widehat{h}_{n}^{2}\right]+Kn_{b}\widetilde{H}_{K}n_{b}^{-2/5}+Kn_{b}\mathbb{E}\left[\left(\widehat{B}-B\right)\widehat{h}_{n}^{2}\right]$
$\leq$ $Kn_{b}\mathbb{E}\left[\alpha^{\prime}(\theta_{0})+B\widehat{h}_{n}^{2}\right]+Kn_{b}O\left(n_{b}^{-2/5}\right)+Kn_{b}\left(\mathbb{E}\left[\widehat{B}-B\right]^{2}\mathbb{E}\left[{\widehat{h}_{n}^{4}}\right]\right)^{1/2}$
$=$ $Kn_{b}\mathbb{E}\left[\alpha^{\prime}(\theta_{0})+B\widehat{h}_{n}^{2}\right]+Kn_{b}O\left(n_{b}^{-2/5}\right)+Kn_{b}O\left(n_{b}^{-1/5}\right)O\left(n^{-1/3}\right),$ (74)
where the last two equalities are due to Corollary 6 and Theorem 2, respectively, and the inequality is because of Cauchy-Schwarz inequality.
Lastly, we study the third term on the RHS of (72).
[…]
$=$ $n_{b}D\left(\frac{4nB^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}{\bm{h^{\top}Ph^{4}}}+o\left(n_{b}^{1/2}\right)O\left(n^{1/6}\right)$
$=$ $D\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}{\bm{c^{\top}Pc^{4}}}n_{b}^{1/2}n^{1/6}+o\left(n_{b}^{1/2}\right)O\left(n^{1/6}\right),$ (75)
where ${\bm{h}^{4}}=[h_{1}^{4},...,h_{K}^{4}]^{\top}$, ${\bm{c}}=[|c_{1}|,...,|c_{K}|]^{\top}$ and ${\bm{c}^{4}}=[c_{1}^{4},...,c_{K}^{4}]^{\top}$ and the second equality is because $\bm{P}$ is a projection matrix. [^7] Specifically, for any $\bm{v}\in\mathcal{C}(\bm{X}_{e})$, $\bm{Pv=0}$. Note that here we omit the lengthy discussion to ensure that the remainder terms have the form $o\left(n_{b}^{1/2}\right)O\left(n^{1/6}\right)$.
Combining (72), (B.2), (B.2) and (B.2) gives
$n\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}=$ $(n_{2}+Kn_{b})\mathbb{E}\left[\alpha^{\prime}(\theta_{0})+B\widehat{h}_{n}^{2}\right]+D\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}{\bm{c^{\top}Pc^{4}}}n_{b}^{1/2}n^{1/6}$
$+n_{2}o\left(n^{-1/3}\right)+Kn_{b}O\left(n_{b}^{-2/5}\right)+Kn_{b}O\left(n_{b}^{-1/5}\right)O\left(n^{-1/3}\right)+o\left(n_{b}^{1/2}\right)O\left(n^{1/6}\right).$
Therefore,
$\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}=$ $\mathbb{E}\left[\alpha^{\prime}(\theta_{0})+B\widehat{h}_{n}^{2}\right]+D\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}{\bm{c^{\top}Pc^{4}}}n_{b}^{1/2}n^{-5/6}+o\left(n^{-1/3}\right)$
$=$ $\alpha^{\prime}(\theta_{0})+\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}n^{-1/3}+o\left(n^{-1/3}\right).$
Proof of (34): The variance of $\widehat{\Delta}_{n_{b},\widehat{h}_{n}}$ is written as
$\mathrm{Var}\left[\widehat{\Delta}_{n_{b},\widehat{h}_{n}}\right]=$ $\frac{1}{n^{2}}\left(\mathrm{Var}\left[\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)\right]+\mathrm{Var}\left[n_{b}\bm{h}^{\top}\bm{P}\bm{\Delta_{h}}\right]+\mathrm{Var}\left[Kn_{b}\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}\right]\right.$
$+2\mathrm{Cov}\left[\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right),n_{b}\bm{h}^{\top}\bm{P}\bm{\Delta_{h}}\right]+2\mathrm{Cov}\left[\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right),Kn_{b}\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}\right]$
$\left.+2\mathrm{Cov}\left[n_{b}\bm{h}^{\top}\bm{P}\bm{\Delta_{h}},Kn_{b}\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}\right]\right).$ (76)
We look at the six terms on the RHS of (B.2) one by one.
Firstly, we consider the first term on the RHS of (B.2).
The first term on the RHS of (B.2)
$=$ $\frac{1}{n^{2}}\mathrm{Var}\left[\mathbb{E}\left[\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)\left|\widehat{h}_{n}\right.\right]\right]+\frac{1}{n^{2}}\mathbb{E}\left[\mathrm{Var}\left[\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)\left|\widehat{h}_{n}\right.\right]\right]$
$=$ $\frac{1}{n^{2}}\mathrm{Var}\left[n_{2}\left(\alpha^{\prime}(\theta_{0})+B\widehat{h}_{n}^{2}+o\left(\widehat{h}_{n}^{2}\right)\right)\right]+\frac{1}{n^{2}}\mathbb{E}\left[n_{2}\frac{\sigma^{2}(\theta_{0})+o(1)}{2\widehat{h}_{n}^{2}}\right]$
$=$ $\frac{n_{2}^{2}}{n^{2}}\mathrm{Var}\left[B\widehat{h}_{n}^{2}+o\left(\widehat{h}_{n}^{2}\right)\right]+\frac{1}{n^{2}}\mathbb{E}\left[n_{2}\frac{\sigma^{2}(\theta_{0})+o(1)}{2\widehat{h}_{n}^{2}}\right]$
$=$ $\frac{n_{2}^{2}o\left(n^{-2/3}\right)}{n^{2}}+\mathbb{E}\left[n_{2}\frac{\sigma^{2}(\theta_{0})+o(1)}{2n^{2}\widehat{h}_{n}^{2}}\right]$
$=$ $\mathbb{E}\left[\frac{n_{2}\sigma^{2}(\theta_{0})}{2n^{2}\widehat{h}_{n}^{2}}\right]+o\left(n^{-2/3}\right),$ (77)
where the second equality is due to (5) and (6) and the fourth equality is because
$n^{2/3}\mathrm{Var}\left[\widehat{h}_{n}^{2}\right]=\mathrm{Var}\left[\left(\frac{\widehat{\bm{\beta}}_{v}}{4\widehat{B}^{2}}\right)^{1/3}\right]=\mathbb{E}\left[\frac{\widehat{\bm{\beta}}_{v}}{4\widehat{B}^{2}}\right]^{2/3}-\mathbb{E}^{2}\left[\frac{\widehat{\bm{\beta}}_{v}}{4\widehat{B}^{2}}\right]^{1/3}\to\frac{\bm{\beta}_{v}^{2/3}}{(4B^{2})^{2/3}}-\left(\frac{\bm{\beta}_{v}^{1/3}}{(4B^{2})^{1/3}}\right)^{2}=0$
implies $\mathrm{Var}\left[\widehat{h}_{n}^{2}\right]=o\left(n^{-2/3}\right)$.
Secondly, we study the second term on the RHS of (B.2).
The second term on the RHS of (B.2)
$=$ $\frac{n_{b}^{2}}{n^{2}}\bm{h}^{\top}\bm{P}\mathrm{Var}[\bm{\Delta_{h}}]\bm{Ph}$
$=$ $\frac{n_{b}^{2}}{n^{2}}\bm{h}^{\top}\bm{P}\left[\begin{array}[]{ccc}\mathrm{Var}\left[\frac{\widehat{\Delta}_{n_{b},h_{1}}}{\left|\widehat{h}_{n}\right|}\right]&...&\mathrm{Cov}\left[\frac{\widehat{\Delta}_{n_{b},h_{1}}}{\left|\widehat{h}_{n}\right|},\frac{\widehat{\Delta}_{n_{b},h_{K}}}{\left|\widehat{h}_{n}\right|}\right]\\ \vdots&&\vdots\\ \mathrm{Cov}\left[\frac{\widehat{\Delta}_{n_{b},h_{K}}}{\left|\widehat{h}_{n}\right|},\frac{\widehat{\Delta}_{n_{b},h_{1}}}{\left|\widehat{h}_{n}\right|}\right]&...&\mathrm{Var}\left[\frac{\widehat{\Delta}_{n_{b},h_{K}}}{\left|\widehat{h}_{n}\right|}\right]\end{array}\right]\bm{Ph}.$
Note that
$\mathrm{Cov}\left[\frac{\widehat{\Delta}_{n_{b},h_{1}}}{\left|\widehat{h}_{n}\right|},\frac{\widehat{\Delta}_{n_{b},h_{K}}}{\left|\widehat{h}_{n}\right|}\right]=\mathrm{Cov}\left[\frac{\widehat{\Delta}_{n_{b},h_{1}}}{h^{*}}+\widehat{\Delta}_{n_{b},h_{1}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right),\frac{\widehat{\Delta}_{n_{b},h_{K}}}{h^{*}}+\widehat{\Delta}_{n_{b},h_{K}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right)\right]$
$=$ $\mathrm{Cov}\left[\frac{\widehat{\Delta}_{n_{b},h_{1}}}{h^{*}},\frac{\widehat{\Delta}_{n_{b},h_{K}}}{h^{*}}\right]+\mathrm{Cov}\left[\frac{\widehat{\Delta}_{n_{b},h_{1}}}{h^{*}},\widehat{\Delta}_{n_{b},h_{K}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right)\right]+\mathrm{Cov}\left[\widehat{\Delta}_{n_{b},h_{1}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right),\frac{\widehat{\Delta}_{n_{b},h_{K}}}{h^{*}}\right]$
$+\mathrm{Cov}\left[\widehat{\Delta}_{n_{b},h_{1}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right),\widehat{\Delta}_{n_{b},h_{K}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right)\right]$
$=$ $\mathrm{Cov}\left[\frac{\widehat{\Delta}_{n_{b},h_{1}}}{h^{*}},\widehat{\Delta}_{n_{b},h_{K}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right)\right]+\mathrm{Cov}\left[\widehat{\Delta}_{n_{b},h_{1}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right),\frac{\widehat{\Delta}_{n_{b},h_{K}}}{h^{*}}\right]$
$+\mathrm{Cov}\left[\widehat{\Delta}_{n_{b},h_{1}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right),\widehat{\Delta}_{n_{b},h_{K}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right)\right]$
$\leq$ $\sqrt{\mathrm{Var}\left[\frac{\widehat{\Delta}_{n_{b},h_{1}}}{h^{*}}\right]\mathrm{Var}\left[\widehat{\Delta}_{n_{b},h_{K}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right)\right]}+\sqrt{\mathrm{Var}\left[\widehat{\Delta}_{n_{b},h_{1}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right)\right]\mathrm{Var}\left[\frac{\widehat{\Delta}_{n_{b},h_{K}}}{h^{*}}\right]}$
$+\sqrt{\mathrm{Var}\left[\widehat{\Delta}_{n_{b},h_{1}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right)\right]\mathrm{Var}\left[\widehat{\Delta}_{n_{b},h_{K}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right)\right]}.$
Because $n^{-1/6}\widehat{\Delta}_{n_{b},h_{k}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right)=o_{p}\left(\widehat{\Delta}_{n_{b},h_{k}}\right)$, for any $k=1,2,...,K$, $\mathrm{Var}\left[\widehat{\Delta}_{n_{b},h_{k}}\left(\frac{1}{\left|\widehat{h}_{n}\right|}-\frac{1}{h^{*}}\right)\right]=n^{1/3}\mathrm{Var}\left[o_{p}\left(\widehat{\Delta}_{n_{b},h_{k}}\right)\right]=o\left(n_{b}^{-4/5}\right)O\left(n^{1/3}\right)$. As a result, $\mathrm{Cov}\left[\frac{\widehat{\Delta}_{n_{b},h_{1}}}{\left|\widehat{h}_{n}\right|},\frac{\widehat{\Delta}_{n_{b},h_{K}}}{\left|\widehat{h}_{n}\right|}\right]=o\left(n_{b}^{-4/5}\right)O\left(n^{1/3}\right)$. Similarly, for any $k=1,2,...,K$,
$\mathrm{Var}\left[\frac{\widehat{\Delta}_{n_{b},h_{k}}}{\left|\widehat{h}_{n}\right|}\right]=$ $\mathrm{Cov}\left[\frac{\widehat{\Delta}_{n_{b},h_{k}}}{\left|\widehat{h}_{n}\right|},\frac{\widehat{\Delta}_{n_{b},h_{k}}}{\left|\widehat{h}_{n}\right|}\right]=\mathrm{Var}\left[\frac{\widehat{\Delta}_{n_{b},h_{k}}}{h^{*}}\right]+o\left(n_{b}^{-4/5}\right)O\left(n^{1/3}\right)$
$=$ $\frac{\sigma^{2}(\theta_{0})}{2n_{b}h_{k}^{2}}\left(\frac{4nB^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/3}+o\left(n_{b}^{-4/5}\right)O\left(n^{1/3}\right).$
Therefore,
The second term on the RHS of (B.2)
$=$ $\frac{n_{b}^{2}}{n^{2}}\bm{h}^{\top}\bm{P}\left[\begin{array}[]{ccc}\mathrm{Var}\left[\frac{\widehat{\Delta}_{n_{b},h_{1}}}{\left|\widehat{h}_{n}\right|}\right]&&\\ &\ddots&\\ &&\mathrm{Var}\left[\frac{\widehat{\Delta}_{n_{b},h_{K}}}{\left|\widehat{h}_{n}\right|}\right]\end{array}\right]\bm{Ph}+\frac{n_{b}^{2}}{n^{2}}n_{b}^{1/10}o\left(n_{b}^{-4/5}\right)O\left(n^{1/3}\right)n_{b}^{1/10}$ (81)
$=$ $\frac{n_{b}}{n^{2}}\left(\frac{nB^{2}\sigma^{4}(\theta_{0})}{2}\right)^{1/3}\bm{h}^{\top}\bm{P}\left[\begin{array}[]{ccc}\frac{1}{h_{1}^{2}}&&\\ &\ddots&\\ &&\frac{1}{h_{K}^{2}}\end{array}\right]\bm{Ph}+o\left(n^{-2/3}\right)$ (85)
$=$ $\frac{n_{b}}{n^{2}}\left(\frac{nB^{2}\sigma^{4}(\theta_{0})}{2}\right)^{1/3}\left|\left|\left[\begin{array}[]{ccc}\frac{1}{|h_{1}|}&&\\ &\ddots&\\ &&\frac{1}{|h_{K}|}\end{array}\right]\bm{Ph}\right|\right|_{2}^{2}+o\left(n^{-2/3}\right)$ (89)
$=$ $\frac{n_{b}}{n^{2}}\left(\frac{nB^{2}\sigma^{4}(\theta_{0})}{2}\right)^{1/3}\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}+o\left(n^{-2/3}\right),$ (90)
where ${\rm{Diag}}(\bm{c^{-1}})={\rm{Diag}}\left(\frac{1}{|c_{1}|},...,\frac{1}{|c_{K}|}\right)$ denotes the diagonal matrix and $||{\bm{v}}||_{2}=\left(\sum_{k=1}^{K}v_{k}^{2}\right)^{1/2}$ for any ${\bm{v}}\in\mathbb{R}^{K}$.
Then, we study the third term on the RHS of (B.2).
The third term on the RHS of (B.2)
$=$ $\frac{K^{2}n_{b}^{2}}{n^{2}}\mathrm{Var}\left[\widehat{\alpha}^{\prime}(\theta_{0})+\widehat{B}\widehat{h}_{n}^{2}\right]$
$=$ $\frac{K^{2}n_{b}^{2}}{n^{2}}\left(\mathrm{Var}\left[\widehat{\alpha}^{\prime}(\theta_{0})\right]+\mathrm{Var}\left[\widehat{B}\widehat{h}_{n}^{2}\right]+2\mathrm{Cov}\left[\widehat{\alpha}^{\prime}(\theta_{0}),\widehat{B}\widehat{h}_{n}^{2}\right]\right)$ (91)
We study the three terms on the RHS of (B.2) one by one. Firstly, according to Corollary 6
[…]
Secondly, we study the second term on the RHS of (B.2).
The second term on the RHS of (B.2)
$=$ $\frac{K^{2}n_{b}^{2}}{n^{2}}\mathrm{Var}\left[B\widehat{h}_{n}^{2}+\left(\widehat{B}-B\right)\widehat{h}_{n}^{2}\right]$
$=$ $\frac{K^{2}n_{b}^{2}}{n^{2}}\left(\mathrm{Var}\left[B\widehat{h}_{n}^{2}\right]+\mathrm{Var}\left[\left(\widehat{B}-B\right)\widehat{h}_{n}^{2}\right]+2\mathrm{Cov}\left[B\widehat{h}_{n}^{2},\left(\widehat{B}-B\right)\widehat{h}_{n}^{2}\right]\right)$
$\leq$ $\frac{K^{2}n_{b}^{2}}{n^{2}}\left(\mathrm{Var}\left[B\widehat{h}_{n}^{2}\right]+\mathbb{E}\left[\left(\widehat{B}-B\right)^{2}\widehat{h}_{n}^{4}\right]+2\mathrm{Cov}\left[B\widehat{h}_{n}^{2},\left(\widehat{B}-B\right)\widehat{h}_{n}^{2}\right]\right)$
$\leq$ $\frac{K^{2}n_{b}^{2}}{n^{2}}\left(\mathrm{Var}\left[B\widehat{h}_{n}^{2}\right]+\left(\mathbb{E}\left[\widehat{B}-B\right]^{4}\mathbb{E}\left[\widehat{h}_{n}^{8}\right]\right)^{1/2}+2\mathrm{Cov}\left[B\widehat{h}_{n}^{2},\left(\widehat{B}-B\right)\widehat{h}_{n}^{2}\right]\right)$
$=$ $\frac{K^{2}n_{b}^{2}}{n^{2}}\left(n^{-2/3}\mathrm{Var}\left[B\left(\frac{\widehat{\bm{\beta}}_{v}}{4\widehat{B}^{2}}\right)^{1/3}\right]+O\left(n_{b}^{-2/5}\right)O\left(n^{-2/3}\right)+2\mathrm{Cov}\left[B\widehat{h}_{n}^{2},\left(\widehat{B}-B\right)\widehat{h}_{n}^{2}\right]\right)$
$=$ $\frac{K^{2}n_{b}^{2}}{n^{2}}\left(o\left(n^{-2/3}\right)+O\left(n_{b}^{-2/5}\right)O\left(n^{-2/3}\right)\right)$
$=$ $O\left(n_{b}^{2}\right)o\left(n^{-8/3}\right)+O\left(n_{b}^{8/5}\right)O\left(n^{-8/3}\right)$
$=$ $o\left(n^{-2/3}\right),$ (93)
where the third equality is because Theorem 2 implies that $\mathbb{E}\left[\widehat{B}-B\right]^{4}=O\left(n_{b}^{-4/5}\right)$ and the fourth equality is because the covariance term is controlled by the two variance terms.
According to Cauchy-Schwarz inequality,
The third term on the RHS of (B.2)
[…]
$=$ $\sqrt{o\left(n^{-2/3}\right)o\left(n^{-2/3}\right)}=o\left(n^{-2/3}\right).$ (94)
Combining (B.2), (92), (B.2) and (B.2) gives that
[…]
Next, we consider the fourth term on the RHS of (B.2).
The fourth term on the RHS of (B.2)
$=$ $\frac{2}{n^{2}}\left\{\mathbb{E}\left[\mathrm{Cov}\left[\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right),n_{b}{\bm{h}^{\top}\bm{P}\bm{\Delta_{h}}}\left|\widehat{h}_{n},{\bm{\Delta_{h}}}\right.\right]\right]\right.$
$+\left.\mathrm{Cov}\left[\mathbb{E}\left[\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)\left|\widehat{h}_{n},{\bm{\Delta_{h}}}\right.\right],\mathbb{E}\left[n_{b}{\bm{h}^{\top}\bm{P}\bm{\Delta_{h}}}\left|\widehat{h}_{n},{\bm{\Delta_{h}}}\right.\right]\right]\right\}$
$=$ $\frac{2}{n^{2}}\mathrm{Cov}\left[n_{2}\left(\alpha^{\prime}(\theta_{0})+B\widehat{h}_{n}^{2}+o_{p}\left(n^{-1/3}\right)\right),n_{b}{\bm{h}^{\top}\bm{P}\bm{\Delta_{h}}}\right]$
$=$ $\frac{2}{n^{2}}\mathrm{Cov}\left[n_{2}\left(B\widehat{h}_{n}^{2}+o_{p}\left(n^{-1/3}\right)\right),n_{b}{\bm{h}^{\top}\bm{P}\bm{\Delta_{h}}}\right]$
$=$ $\frac{2n_{2}n_{b}}{n^{2}}B{\bm{h}^{\top}\bm{P}}\mathrm{Cov}\left[\widehat{h}_{n}^{2},\bm{\Delta_{h}}\right]+\mbox{remainder terms}$
$=$ $\frac{2n_{2}n_{b}^{9/10}}{n^{2}}B{\bm{c}^{\top}\bm{P}}\mathrm{Cov}\left[\widehat{h}_{n}^{2},\bm{\Delta_{h}}\right]+\mbox{remainder terms},$
where the first equality is due to the law of total covariance. The second equality is because when ${\bm{\Delta_{h}}}$ is given, $n_{b}{\bm{h}^{\top}\bm{P}\bm{\Delta_{h}}}$ is a constant and $\mathrm{Cov}\left[\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right),n_{b}{\bm{h}^{\top}\bm{P}\bm{\Delta_{h}}}\right]=0$. According to the Cauchy-Schwarz inequality,
$\mathrm{Cov}\left[\widehat{h}_{n}^{2},\frac{\widehat{\Delta}_{n_{b},h_{1}}}{\left|\widehat{h}_{n}\right|}\right]\leq\sqrt{\mathrm{Var}\left[\widehat{h}_{n}^{2}\right]\mathrm{Var}\left[\frac{\widehat{\Delta}_{n_{b},h_{1}}}{\left|\widehat{h}_{n}\right|}\right]}=\sqrt{o\left(n^{-2/3}\right)O\left(n_{b}^{-4/5}n^{1/3}\right)}=o\left(n^{-1/6}\right)O\left(n_{b}^{-2/5}\right).$
Therefore,
[…]
By Cauchy-Schwarz inequality, we have
The fifth term on the RHS of (B.2)
[…]
$=$ $o\left(n^{-2/3}\right)$ (97)
and
The sixth term on the RHS of (B.2)
[…]
$=$ $o\left(n^{-2/3}\right).$ (98)
Combining (B.2), (B.2), (90), (95), (96), (B.2) and (B.2) gives
$\mathrm{Var}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]=$ $\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2n^{2}}\right)^{1/3}+\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2}\right)^{1/3}\frac{n_{b}}{n}\left[\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}-K\right]n^{-2/3}+o\left(n^{-2/3}\right).$
The proof is complete.
B.3 Remarks on Theorem 4 and Proposition 5
Denote $\mathcal{C}(\bm{X}_{e})$ the column space of $\bm{X}_{e}$. Then $\bm{P}$ is a projection matrix onto $\mathcal{C}(\bm{X}_{e})^{\bot}$ which is the orthogonal complement space of $\mathcal{C}(\bm{X}_{e})$, i.e., $\mathcal{C}(\bm{X}_{e})\oplus\mathcal{C}(\bm{X}_{e})^{\bot}=\mathbb{R}^{K}$. Therefore, for any ${\bm{v}}\in\mathbb{R}^{K}$, we can decompose it as ${\bm{v}}={\bm{v}}_{1}+{\bm{v}}_{2}$, where ${\bm{v}}_{1}\in\mathcal{C}(\bm{X}_{e})$ and ${\bm{v}}_{2}\in\mathcal{C}(\bm{X}_{e})^{\bot}$. Subsequently, applying $\bm{P}$ to ${\bm{v}}$ results in $\bm{P}{\bm{v}}={\bm{v}}_{2}$. If the angle between ${\bm{v}}$ and $\mathcal{C}(\bm{X}_{e})$ is small, $||\bm{P}{\bm{v}}||_{2}$, which is equal to $||{\bm{v}}||_{2}$, is also small.
According to the property of projection matrix and Cauchy-Schwarz inequality, $|{\bm{c^{\top}Pc^{4}}}|=|{\bm{c^{\top}P^{\top}Pc^{4}}}|\leq||{\bm{Pc}}||_{2}||{\bm{Pc^{4}}}||_{2}$. If the angles between $\bm{c}$, $\bm{c^{4}}$ and $\mathcal{C}(\bm{X}_{e})$ are small, which are typically the case, both $||{\bm{Pc}}||_{2}$ and $||{\bm{Pc^{4}}}||_{2}$ are small, then $|{\bm{c^{\top}Pc^{4}}}|$ is manageable.
For example, if ${\bm{c}}=[1,1.5,2,...,5.5]^{\top}$, then ${\bm{c}^{4}}=[1,1.5^{4},2^{4},...,5.5^{4}]^{\top}$. Denote the angles between ${\bm{c}}$, ${\bm{c}^{4}}$ and $\mathcal{C}({\bm{X}})$ by $\vartheta$ and $\widetilde{\vartheta}$, respectively. Then, we can easily calculate that $\cos\vartheta\geq 0.9693$ and $\cos\widetilde{\vartheta}\geq 0.9595$, both of which are close to 1. In other words, both $\vartheta$ and $\widetilde{\vartheta}$ are close to 0.
B.4 Proof of Proposition 5
It is obvious that (36) holds when $n_{b}=o(n)$. In the following, we consider the bias and variance terms of case 2 (i.e., $\underset{n_{b}\to\infty}{\lim}\underset{n\to\infty}{\lim}Kn_{b}/n=r$, where $0<r\leq 1$) in Proposition 5 separately.
  • $\mathbb{E}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]-\alpha^{\prime}(\theta_{0})$:
    • i.
      If we have
      $-1-{\rm{sign}}(B)\leq\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}D{\bm{c^{\top}Pc^{4}}}\leq 1-{\rm{sign}}(B),$
      then,
      $\left(\mathbb{E}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]-\alpha^{\prime}(\theta_{0})\right)^{2}$
      $=$ $\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{2/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/3}\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}\left(\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}+{\rm{sign}}(B)\sigma(\theta_{0})\right)$
      $=$ $\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{2/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/3}\sqrt{\frac{r}{K}}D{\bm{c^{\top}Pc^{4}}}\left(\sqrt{\frac{r}{K}}D{\bm{c^{\top}Pc^{4}}}+{\rm{sign}}(B)\sigma(\theta_{0})\right)$
      $\leq$ $\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{2/3}.$
      Therefore,
      $\left|\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}-\alpha^{\prime}(\theta_{0})\right|\leq\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/6}+o\left(n^{-1/3}\right).$
    • ii.
      Else, from (33) in Theorem 4, we have
      $\left|\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}-\alpha^{\prime}(\theta_{0})\right|=$ $\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}n^{-1/3}+o\left(n^{-1/3}\right)$
      $\leq$ $\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/6}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}\sqrt{\frac{r}{K}}\left|D{\bm{c^{\top}Pc^{4}}}\right|n^{-1/3}+o\left(n^{-1/3}\right)$
      $\leq$ $\left(1+\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}\left|D{\bm{c^{\top}Pc^{4}}}\right|\right)\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/6}+o\left(n^{-1/3}\right).$
  • $\mathrm{Var}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]$: From Theorem 4, it suffices to prove $\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}\leq K$.
    $\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}\leq\frac{1}{\min_{k\in\{1,...,K\}}c_{k}^{2}}\left|\left|\bm{Pc}\right|\right|_{2}^{2}=\frac{\min_{x,y}\left|\left|x\bm{1}+y\bm{c^{2}}-\bm{c}\right|\right|_{2}^{2}}{\min_{k\in\{1,...,K\}}c_{k}^{2}}$
    $\leq$ $\frac{K\left|\left|\bm{c}\right|\right|_{2}^{2}-\left|\left|\bm{c}\right|\right|_{1}^{2}}{K\min_{k\in\{1,...,K\}}c_{k}^{2}}\leq\frac{K\max_{k\in\{1,...,K\}}c_{k}^{2}-K\min_{k\in\{1,...,K\}}c_{k}^{2}}{\min_{k\in\{1,...,K\}}c_{k}^{2}}\leq K,$
    where the first inequality is because that for any matrix ${\bm{A}}$ and vector ${\bm{v}}$, $||{\bm{A}v}||_{2}\leq||{\bm{A}}||_{2}||{\bm{v}}||_{2}$, the first equality is due to the property of projection matrix, the second inequality holds by taking $x=\left|\left|\bm{c}\right|\right|_{1}/K$ and $y=0$ and the last inequality is attributed to $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$.
The proof is complete.
B.5 A similar conclusion to Theorem 4 for explaining the variance reduction and bias reduction in Remark 2
To illustrate that variance reduction and possible bias reduction occur when substituting $\mathbb{E}[\Delta(h_{k})],k=1,...,K$ with their estimators, we refrain from using the substitution and instead assume that $\mathbb{E}\Delta(h_{k}),k=1,...,K$ are all precisely known. Then, the new DSR-CFD estimator, denoted by $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$, is also formulated as (32) but with $\Delta_{j}^{t}(h_{k})$ replaced by
$\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\Delta_{j}(h_{k})-\mathbb{E}\Delta(h_{k})\right)+\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}.$
In the following, we provide the asymptotic properties of $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$.
Proposition 9.
Under the same conditions as those of Theorem 4, we have
$\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}=\alpha^{\prime}(\theta_{0})+\left(\dfrac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+o\left(n^{-1/3}\right),$ (99)
$\mathrm{Var}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}\right]=\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{2n^{2}}\right)^{1/3}+o\left(n^{-2/3}\right).$ (100)
Proof
We prove (99) and (100) separately.
Proof of (99): It follows that
$\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}=\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+n_{b}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\widehat{\Delta}_{n_{b},h_{k}}-\mathbb{E}\Delta(h_{k})\right)+Kn_{b}\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}\right).$ (101)
According to Appendix B.2, to prove (99), it suffices to prove the expectation of the second term on the RHS of (101) is $o\left(n^{-1/3}\right)$. The proof is as follows.
$\mathbb{E}\left[\frac{n_{b}}{n}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\widehat{\Delta}_{n_{b},h_{k}}-\mathbb{E}\Delta(h_{k})\right)\right]$
$=$ $\frac{n_{b}}{n}\sum_{k=1}^{K}|h_{k}|\mathbb{E}\left[\frac{\widehat{\Delta}_{n_{b},h_{k}}-\mathbb{E}\Delta(h_{k})}{\left|\widehat{h}_{n}\right|}\right]$
$=$ $\frac{n_{b}}{n}\sum_{k=1}^{K}|h_{k}|\left(\mathbb{E}\left[\frac{\widehat{\Delta}_{n_{b},h_{k}}}{\left|\widehat{h}_{n}\right|}\right]-\mathbb{E}\Delta(h_{k})\mathbb{E}\left[\frac{1}{\left|\widehat{h}_{n}\right|}\right]\right)$
$=$ $\frac{n_{b}}{n}\left(\frac{4nB^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}\sum_{k=1}^{K}|h_{k}|\left(\alpha^{\prime}(\theta_{0})+Bh_{k}^{2}+Dh_{k}^{4}-\mathbb{E}\Delta(h_{k})\right)+o\left(n_{b}^{1/2}\right)O\left(n^{-5/6}\right)$
$=$ $o\left(n_{b}^{1/2}\right)O\left(n^{-5/6}\right)=o\left(n^{-1/3}\right).$
Proof of (100): Note that
$\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}=\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+n_{b}\bm{h}^{\top}\bm{I}\bm{\Delta_{h}}+Kn_{b}\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}-n_{b}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\mathbb{E}\Delta(h_{k})\right).$ (102)
It is straightforward to show that the variance of the last term on the RHS of (102) is $o\left(n^{-2/3}\right)$, so the essence is calculating the variance of the first three terms on the RHS of (102). Compared to $\widehat{\Delta}_{n,\widehat{h}_{n}}$ in Appendix B.2, the first three terms of (102) simply replace the matrix $\bm{P}$ with the identity matrix $\bm{I}$. Therefore, following the technique used in Appendix B.2, we get
$\mathrm{Var}\left[\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+n_{b}\bm{h}^{\top}\bm{I}\bm{\Delta_{h}}+Kn_{b}\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}\right)\right]$
$=$ $\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2n^{2}}\right)^{1/3}+\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2}\right)^{1/3}\frac{n_{b}}{n}\left[\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Ic}\right|\right|_{2}^{2}-K\right]n^{-2/3}+o\left(n^{-2/3}\right)$
$=$ $\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2n^{2}}\right)^{1/3}+o\left(n^{-2/3}\right).$
Using the Cauchy-Schwarz inequality, we have
$\mathrm{Var}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}\right]=\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{2n^{2}}\right)^{1/3}+o\left(n^{-2/3}\right).$
The proof is complete.
Appendix C C Additional Experiment Results
Table 4: Comparison of the bias of the DSR-CFD and Opt-CFD estimators in Example 2 when $x=0.4$.
[…]
Budget $=10^{6}$ $-0.0090$ $-0.0063$ 1.4286
Budget $=2\times 10^{6}$ $-0.0067$ $-0.0051$ 1.3137
Budget $=4\times 10^{6}$ $-0.0051$ $-0.0039$ 1.3077
Budget $=8\times 10^{6}$ $-0.0040$ $-0.0031$ 1.2903